### v5

# Local Topological Constraints on Berry Curvature in Spin–Orbit Coupled BECs

Abstract
We establish a local topological obstruction to the simultaneous flattening of Berry curvature in spin–orbit-coupled Bose–Einstein condensates (SOC BECs), which remains valid even when the global Chern number vanishes. For a generic two-component SOC BEC, the extended parameter space is the total space $M$ of a principal $U(1)_{+}\times U(1)_{-}$ bundle over the Brillouin torus $T^{2}_{\mathrm{BZ}}$. On $M$, we construct a Kaluza–Klein metric and a natural metric connection $\nabla^{C}$ whose torsion 3-form encodes the synthetic gauge fields.
Under the physically relevant assumption of constant Berry curvatures, the harmonic part of this torsion defines a mixed cohomology class
$\omega]\in\bigl(H^{2}(T^{2}_{\mathrm{BZ}})\otimes H^{1}(S^{1}_{\phi_{+}})\bigr)\oplus\bigl(H^{2}(T^{2}_{\mathrm{BZ}})\otimes H^{1}(S^{1}_{\phi_{-}})\bigr)$
with mixed tensor rank $r=1$. By adapting the Pigazzini–Toda (PT) lower bound to the Kaluza–Klein setting through explicit pointwise curvature analysis, we demonstrate that the obstruction kernel $\mathcal{K}$ vanishes for the physical metric, yielding the sharp inequality $\dim\mathfrak{hol}^{\mathrm{off}}(\nabla^{C})\geq 1$. This bound forces the existence of at least one off-diagonal curvature operator, preventing the complete gauging-away of Berry phases even in regimes with zero net topological charge. This work provides the first cohomological lower bound, based on the PT framework, certifying locally irremovable curvature in SOC BECs beyond the Chern-number paradigm.
Keywords: Berry curvature; spin–orbit-coupled Bose–Einstein condensates; synthetic gauge fields; holonomy with torsion; mixed cohomology; Kaluza–Klein geometry.
MSC (2020): 53C29, 53C80, 81V45, 82B10, 51P05, 55S20, 57R19, 57R56, 58A12.
1 Introduction
The study of spin–orbit-coupled (SOC) Bose–Einstein condensates (BECs) stands at a rich confluence of differential geometry, quantum many-body physics, and synthetic gauge theory. This interdisciplinary effort is part of the broader pursuit of artificial gauge fields in engineered quantum systems, spanning ultracold atoms, photonics, and solid-state platforms [1]. Advances in generating artificial gauge fields with Raman lasers have made these systems a versatile laboratory for exploring Berry curvature, synthetic magnetic fields, and non-Abelian textures with unprecedented control [6, 7, 14]. Experimental realizations in laboratories such as NIST [9] have demonstrated the ability to engineer synthetic spin-orbit coupling in ultracold atomic gases, providing a controllable platform for investigating topological phenomena in quantum matter.
A persistent theoretical challenge is to identify local topological obstructions that constrain the Berry curvature even when global invariants—such as the first Chern number—vanish. While the Chern number $c_{1}=\frac{1}{2\pi}\int_{T^{2}_{\mathrm{BZ}}}F$ quantizes the net flux through the Brillouin zone, its vanishing does not guarantee that the curvature $F$ can be locally flattened everywhere; geometric obstructions may prevent the simultaneous annihilation of curvature components along independent directions in the parameter space. Understanding such local constraints is crucial for assessing the robustness of topological features in quantum systems and for designing protocols that exploit geometry beyond global topological charges.
This work addresses that challenge by bringing a recent geometric result into physical focus. In [12], a lower bound was established for the off-diagonal holonomy of metric connections with totally skew-symmetric torsion on product manifolds. The bound is expressed in terms of a mixed de Rham cohomology class $\omega]\in H^{p}(M_{1})\otimes H^{q}(M_{2})$ associated with the harmonic part of the torsion 3-form:
$\dim\mathfrak{hol}^{\mathrm{off}}(\nabla^{C})\geq r^{\sharp}=\min(r-\dim\mathcal{K},1),$ (1.1)
where $r=\mathrm{rank}_{\mathbb{R}}([\omega]_{\mathrm{mixed}})$ is the minimal number of simple tensors needed to represent the mixed cohomology class, $\mathcal{K}$ is an obstruction kernel encoding the intersection of this class with spaces of parallel forms, and $\mathfrak{hol}^{\mathrm{off}}(\nabla^{C})$ denotes the off-diagonal part of the holonomy algebra that mixes the two factor manifolds. This bound is a topological invariant that persists under metric deformations preserving the parallel-form strata.
Methodological adaptation: From product to Kaluza–Klein geometry
While Theorem 5.2 of [12] is formally stated for Riemannian product metrics, the physical metric $g_{M}$ in our SOC BEC model is a Kaluza–Klein metric induced by the synthetic gauge fields $A^{(\pm)}$. We do not apply the theorem as a black box. Instead, we extend its reach to the present geometry through a three-step strategy:
  1. (i)
    Deformation. We construct a smooth one-parameter family of metrics $\{g_{\varepsilon}\}_{\varepsilon\in[0,1]}$ that interpolates between a Riemannian product metric $g_{0}=g_{\mathrm{BZ}}\oplus d\phi_{+}^{2}\oplus d\phi_{-}^{2}$ at $\varepsilon=0$ and the physical Kaluza–Klein metric $g_{M}=g_{1}$ at $\varepsilon=1$ (Section 4). This deformation is achieved by continuously introducing the gauge potentials $A^{(\pm)}$ into the fiber metric via the connection 1-forms $\Theta^{(\pm)}_{\varepsilon}=d\phi_{\pm}+\varepsilon\pi^{*}A^{(\pm)}$.
  2. (ii)
    Direct verification for $\varepsilon>0$. For every $\varepsilon>0$, we prove that the obstruction kernel $\mathcal{K}_{\varepsilon}$ vanishes (Theorem 3.14), yielding a non-trivial reduced rank $r^{\sharp}_{\varepsilon}=1$. By explicit computation in an adapted gauge (Appendix B), we evaluate the curvature $R^{C_{\varepsilon}}(X,Z)$ for mixed inputs $X\in\mathcal{H}_{p}$, $Z\in\mathcal{V}_{p}$ and show that its off-diagonal projection $\pi_{\mathrm{off}}(R^{C_{\varepsilon}}(X,Z))\neq 0$ at each point $p\in M$, demonstrating that at least one linearly independent off-diagonal curvature operator exists in the holonomy algebra. This establishes the bound $\dim\mathfrak{hol}^{\mathrm{off}}(\nabla^{C_{\varepsilon}})\geq 1$ for all $\varepsilon\in(0,1]$ through a direct analysis of the curvature tensor, without relying on continuity arguments.
  3. (iii)
    Application to the physical metric. Since the bound holds for every $\varepsilon\in(0,1]$ and the physical connection corresponds to $\varepsilon=1$, we obtain the desired inequality for the physical system. The triviality of the bound at $\varepsilon=0$ (where $r^{\sharp}_{0}=0$ due to complete absorption by parallel forms in the product geometry) highlights the essential role of the Kaluza–Klein coupling in generating irreducible topological constraints.
We take as the extended parameter space the smooth manifold
$M=T^{2}_{\mathrm{BZ}}\times S^{1}_{\phi_{+}}\times S^{1}_{\phi_{-}},$
where $T^{2}_{\mathrm{BZ}}$ is the Brilloin torus parametrized by crystal momenta $(k_{x},k_{y})$, and $S^{1}_{\phi_{+}}$, $S^{1}_{\phi_{-}}$ represent the global U(1) phase (associated with particle-number conservation) and the relative U(1) phase between the two spin components, respectively. Our analysis shows that under the physically motivated assumption of constant Berry curvatures $F^{(\pm)}=c^{(\pm)}\mathrm{vol}_{\mathrm{BZ}}$ (Assumption 2.3), the mixed tensor rank is $r=1$. The vanishing of the obstruction kernel for the Kaluza–Klein geometry (Theorem 2.13) then yields the sharp inequality:
$\dim\mathfrak{hol}^{\mathrm{off}}(\nabla^{C})\geq 1\quad\text{for every point }p\in M.$ (1.2)
This lower bound implies that momentum degrees of freedom are fundamentally “locked” to the phase sectors in a topologically robust manner: at least one independent off-diagonal curvature operator persists, mixing the Brillouin zone with the phase directions even when the total Chern number $c^{(+)}_{1}+c^{(-)}_{1}$ vanishes. Such mixing cannot be removed by any smooth gauge transformation or metric deformation that preserves the cohomology classes $F^{(\pm)}]$. This result reveals a local topological obstruction invisible to the Chern number and suggests new interferometric protocols to detect these geometric constraints through measurements of correlated Berry phases in the two U(1) sectors.
Organization of the paper
The paper is organized as follows. Section 2 constructs the Kaluza–Klein metric and defines the metric connection with torsion encoding the synthetic gauge fields. Section 3 analyzes the mixed cohomology structure, computes the tensor rank, and determines the obstruction kernel for both the product and deformed metrics. Section 4 presents the deformation family and establishes the non-trivial lower bound through direct curvature analysis. Section 5 interprets the bound in the context of SOC BECs and discusses its physical consequences. Section 7 provides illustrative examples, including the paradigmatic case of vanishing total Chern flux. Finally, Section 9 summarizes the results and outlines directions for future work. Technical details on the Levi-Civita connection and the explicit curvature computation are relegated to Appendices A and B, respectively.
2 Geometric Construction of the Extended Parameter Space
A generic two-component spin–orbit-coupled Bose–Einstein condensate (SOC BEC) confined in a quasi-two-dimensional toroidal trap is characterized by three independent, experimentally accessible degrees of freedom: the crystal momenta $(k_{x},k_{y})$ spanning a Brillouin zone, a global $U(1)$ phase $\phi_{+}$ associated with particle-number conservation, and a relative $U(1)$ phase $\phi_{-}$ between the two internal spin states [9, 11].
The natural mathematical framework for this system is a principal bundle with structure group $U(1)_{+}\times U(1)_{-}$, whose total space $M$ serves as the extended parameter space. Topologically, $M$ is diffeomorphic to the four-dimensional torus $T^{4}$, but its geometric structure is that of a Kaluza–Klein manifold when equipped with the synthetic gauge fields.
2.1 Global geometric structure and the product approximation
For the purpose of applying the PT lower bound, we work in a framework where the parameter space admits a global product structure, possibly up to a smooth deformation. Although the parameter space $M=T^{2}_{\mathrm{BZ}}\times S^{1}_{\phi_{+}}\times S^{1}_{\phi_{-}}$ is a global topological product, the Berry curvatures $F^{(\pm)}$ are associated to the quantum eigenbundles over $T^{2}_{\mathrm{BZ}}$, rather than to line bundles over $M$. Consequently, non-vanishing Chern classes $c_{1}^{(\pm)}\neq 0$ are fully compatible with the product structure of $M$.
Definition 2.1 (Extended parameter space as a product manifold).
We consider the smooth manifold
$M=T^{2}_{\mathrm{BZ}}\times S^{1}_{\phi_{+}}\times S^{1}_{\phi_{-}}\cong T^{4},$
equipped with a family of Riemannian metrics $\{g_{\varepsilon}\}_{\varepsilon\in[0,1]}$ such that:
  1. 1.
    For $\varepsilon=0$, $g_{0}=g_{\mathrm{BZ}}\oplus d\phi_{+}^{2}\oplus d\phi_{-}^{2}$ is a Riemannian product metric.
  2. 2.
    For $\varepsilon=1$, $g_{1}=g_{M}$ is the physical Kaluza–Klein metric given by (2.1).
  3. 3.
    The family is smooth in $\varepsilon$ and preserves the orthogonal splitting $TM=\mathcal{H}_{\varepsilon}\oplus\mathcal{V}$, where $\mathcal{V}=\ker d\pi$ is the vertical distribution (independent of $\varepsilon$) and $\mathcal{H}_{\varepsilon}$ is the $g_{\varepsilon}$-orthogonal complement.
Remark 2.2 (Topological vs. geometric product structure).
While the underlying smooth manifold is a product, the physical metric $g_{M}$ is not a Riemannian product unless the connections $A^{(\pm)}$ vanish. The family $\{g_{\varepsilon}\}$ interpolates between the geometric product structure ($\varepsilon=0$) and the physical Kaluza–Klein structure ($\varepsilon=1$). This deformation is essential to apply the PT theorem, which requires a genuine Riemannian product structure.
2.2 Constant Berry curvature assumption
To make the geometric analysis fully explicit and to cover a wide range of experimentally relevant configurations, we work under the following assumption:
Assumption 2.3 (Constant Berry curvatures).
The Berry curvature 2-forms are constant multiples of the Brillouin-zone volume form:
$F^{(\pm)}=dA^{(\pm)}=c^{(\pm)}\operatorname{vol}_{\mathrm{BZ}},\qquad c^{(\pm)}\in\mathbb{R}\setminus\{0\},$
where $\operatorname{vol}_{\mathrm{BZ}}=dk_{x}\wedge dk_{y}$. The constants $c^{(\pm)}$ are related to the Chern numbers by
$c_{1}^{(\pm)}=\frac{1}{2\pi}\int_{T^{2}_{\mathrm{BZ}}}F^{(\pm)}=\frac{\operatorname{Area}(T^{2}_{\mathrm{BZ}})}{2\pi}\,c^{(\pm)}.$
This condition is satisfied in numerous SOC BEC realizations, including linear Rashba–Dresselhaus couplings and uniform synthetic magnetic fields.
Remark 2.4 (Topological invariance).
The cohomology classes $F^{(\pm)}]\in H^{2}(T^{2}_{\mathrm{BZ}};\mathbb{R})$ are topological invariants. Under Assumption 2.3, these classes are proportional: $F^{(-)}]=\lambda[F^{(+)}]$ with $\lambda=c^{(-)}/c^{(+)}$. Any deformation of the connection within its cohomology class preserves these invariants, ensuring the robustness of our results.
2.3 The Kaluza–Klein metric and its deformation to a product
We construct a family of Riemannian metrics on $M$ that interpolates between the physical Kaluza–Klein metric and a product metric. Let $g_{\mathrm{BZ}}=dk_{x}^{2}+dk_{y}^{2}$ be the flat metric on the base torus $T^{2}_{\mathrm{BZ}}$.
Definition 2.5 (Deformation family of Kaluza–Klein metrics).
For $\varepsilon\in[0,1]$, define the metric on $M$ by:
$g_{\varepsilon}=\pi^{*}g_{\mathrm{BZ}}+\bigl(d\phi_{+}+\varepsilon\,\pi^{*}A^{(+)}\bigr)^{2}+\bigl(d\phi_{-}+\varepsilon\,\pi^{*}A^{(-)}\bigr)^{2},$ (2.1)
where $\pi:M\to T^{2}_{\mathrm{BZ}}$ is the projection onto the first factor. For $\varepsilon=0$ we obtain the product metric $g_{0}=g_{\mathrm{BZ}}\oplus d\phi_{+}^{2}\oplus d\phi_{-}^{2}$, and for $\varepsilon=1$ we obtain the physical Kaluza–Klein metric $g_{1}=g_{M}$.
Remark 2.6 (Global well-definedness of the metric family).
The metrics $g_{\varepsilon}$ are globally well-defined on the total space $M$ of the principal $U(1)_{+}\times U(1)_{-}$-bundle over $T_{\mathrm{BZ}}^{2}$. The 1-forms $A^{(\pm)}$ are connection 1-forms on local trivializations; the quantities $(d\phi_{\pm}+\varepsilon\pi^{*}A^{(\pm)})^{2}$ are invariant under gauge transformations $A^{(\pm)}\to A^{(\pm)}+d\lambda^{(\pm)}$ because the fiber coordinates $\phi_{\pm}$ transform as $\phi_{\pm}\to\phi_{\pm}-\varepsilon\lambda^{(\pm)}$. Consequently, the metric $g_{\varepsilon}$ descends to a well-defined Riemannian metric on $M$ even when the curvatures $F^{(\pm)}$ have non-zero Chern classes.
Theorem 2.7 (Properties of the deformation family).
The family of metrics $\{g_{\varepsilon}\}_{\varepsilon\in[0,1]}$ defined on $M$ satisfies the following properties:
  1. (i)
    Regularity: Each $g_{\varepsilon}$ is a smooth, positive-definite Riemannian metric on $M$. The projection $\pi:(M,g_{\varepsilon})\to(T^{2}_{\mathrm{BZ}},g_{\mathrm{BZ}})$ is a Riemannian submersion for every $\varepsilon\in[0,1]$.
  2. (ii)
    Orthogonality: The vertical distribution $\mathcal{V}=\ker d\pi=\operatorname{span}\{\partial_{\phi_{+}},\partial_{\phi_{-}}\}$ is $g_{\varepsilon}$-orthogonal to the horizontal distribution $\mathcal{H}_{\varepsilon}$, where the latter is defined as the kernel of the connection 1-forms $\Theta^{(\pm)}_{\varepsilon}=d\phi_{\pm}+\varepsilon\pi^{*}A^{(\pm)}$.
  3. (iii)
    Fiber Isotropy: Each fibre $\pi^{-1}(k)\cong S^{1}\times S^{1}$ inherits a flat metric $g_{\text{fibre}}=d\phi_{+}^{2}+d\phi_{-}^{2}$ which is independent of both the base point $k\in T^{2}_{\mathrm{BZ}}$ and the deformation parameter $\varepsilon$.
  4. (iv)
    Geodesic Geometry: The fibres are totally geodesic submanifolds of $(M,g_{\varepsilon})$ for every $\varepsilon$, and the horizontal distribution $\mathcal{H}_{\varepsilon}$ has a curvature proportional to $\varepsilon F^{(\pm)}$.
Proof.
(i)-(ii) By construction, $g_{\varepsilon}=\pi^{*}g_{\mathrm{BZ}}+(\Theta^{(+)}_{\varepsilon})^{2}+(\Theta^{(-)}_{\varepsilon})^{2}$. Since $g_{\mathrm{BZ}}$ is positive definite and the vertical forms are linearly independent, $g_{\varepsilon}$ is a valid Riemannian metric. The orthogonality follows from the fact that $g_{\varepsilon}(X,V)=0$ whenever $X\in\mathcal{H}_{\varepsilon}$ (so $\Theta^{(\pm)}_{\varepsilon}(X)=0$) and $V\in\mathcal{V}$ (so $d\pi(V)=0$).
(iii) Restricting $g_{\varepsilon}$ to the vertical distribution $\mathcal{V}$ means evaluating it on vectors $V$ such that $d\pi(V)=0$. In this case, $\pi^{*}A^{(\pm)}(V)=A^{(\pm)}(d\pi(V))=0$, so $\Theta^{(\pm)}_{\varepsilon}|_{\mathcal{V}}=d\phi_{\pm}$. Thus $g_{\varepsilon}|_{\mathcal{V}}=d\phi_{+}^{2}+d\phi_{-}^{2}$, which is constant and independent of $\varepsilon$.
(iv) The fibers are flat tori, and the connection potentials $A^{(\pm)}$ depend only on the base coordinates. The vanishing of the second fundamental form of the fibers follows from the Kaluza-Klein structure with constant fiber metrics. The curvature of the horizontal distribution is given by $d\Theta^{(\pm)}_{\varepsilon}=\varepsilon dA^{(\pm)}=\varepsilon F^{(\pm)}$, confirming that $\varepsilon$ scales the synthetic magnetic field without altering the fiber geometry. ∎
2.4 Metric connection with totally skew-symmetric torsion
For each $\varepsilon\in[0,1]$, we define a metric connection $\nabla^{C_{\varepsilon}}$ with torsion that encodes the physical Berry curvature.
Definition 2.8 (Torsion 3-form family).
For each $\varepsilon$, define the torsion 3-form:
$T_{\varepsilon}:=F^{(+)}\wedge(d\phi_{+}+\varepsilon\pi^{*}A^{(+)})+F^{(-)}\wedge(d\phi_{-}+\varepsilon\pi^{*}A^{(-)})\in\Omega^{3}(M).$
Under Assumption 2.3, this becomes:
$T_{\varepsilon}=c^{(+)}\operatorname{vol}_{\mathrm{BZ}}\wedge(d\phi_{+}+\varepsilon\pi^{*}A^{(+)})+c^{(-)}\operatorname{vol}_{\mathrm{BZ}}\wedge(d\phi_{-}+\varepsilon\pi^{*}A^{(-)}).$
Lemma 2.9 (Cohomological invariance of the torsion class).
For all $\varepsilon\in[0,1]$, the de Rham cohomology class $T_{\varepsilon}]\in H^{3}(M;\mathbb{R})$ is constant and equal to
$T_{\varepsilon}]=[F^{(+)}]\otimes[d\phi_{+}]+[F^{(-)}]\otimes[d\phi_{-}]\quad\text{under the Künneth isomorphism.}$
Proof.
We verify that $T_{\varepsilon}$ is closed for every $\varepsilon\in[0,1]$. Using $dF^{(\pm)}=0$ (curvatures are closed) and $d(d\phi_{\pm})=0$, we compute:
$dT_{\varepsilon}=-F^{(+)}\wedge d(\varepsilon\pi^{*}A^{(+)})-F^{(-)}\wedge d(\varepsilon\pi^{*}A^{(-)})=-\varepsilon\left(F^{(+)}\wedge F^{(+)}+F^{(-)}\wedge F^{(-)}\right).$
Under Assumption 2.3, $F^{(\pm)}=c^{(\pm)}dk_{x}\wedge dk_{y}$, hence
$F^{(\pm)}\wedge F^{(\pm)}=c^{(\pm)2}(dk_{x}\wedge dk_{y})\wedge(dk_{x}\wedge dk_{y})=0$
by antisymmetry. Thus $dT_{\varepsilon}=0$ for all $\varepsilon$. Under the Künneth decomposition
$H^{3}(M;\mathbb{R})\cong\left(H^{2}(T^{2}_{\mathrm{BZ}})\otimes H^{1}(S^{1}_{\phi_{+}}\times S^{1}_{\phi_{-}})\right)\oplus\left(H^{1}(T^{2}_{\mathrm{BZ}})\otimes H^{2}(S^{1}_{\phi_{+}}\times S^{1}_{\phi_{-}})\right),$
the class $T_{0}]=[F^{(+)}\wedge d\phi_{+}+F^{(-)}\wedge d\phi_{-}]$ decomposes as
$T_{0}]=[F^{(+)}]\otimes[d\phi_{+}]+[F^{(-)}]\otimes[d\phi_{-}]$
in the $(2,1)$-component (the $(1,2)$-component vanishes since $H^{2}(S^{1}\times S^{1})=0$). For $T_{\varepsilon}=T_{0}+\varepsilon(F^{(+)}\wedge\pi^{*}A^{(+)}+F^{(-)}\wedge\pi^{*}A^{(-)})$, the additional terms are 3-forms of bidegree $(3,0)$ in the Künneth decomposition, being wedge products of forms pulled back from $T^{2}_{\mathrm{BZ}}$. Since
$H^{3}(T^{2}_{\mathrm{BZ}})=0\quad(\text{as }\dim T^{2}_{\mathrm{BZ}}=2<3),$
these terms contribute zero to the cohomology class. Therefore
$T_{\varepsilon}]=[F^{(+)}]\otimes[d\phi_{+}]+[F^{(-)}]\otimes[d\phi_{-}]=[T_{0}]$
for all $\varepsilon\in[0,1]$. ∎
Definition 2.10 (Metric connection family).
For each $\varepsilon\in[0,1]$, define the $(2,1)$-tensor $K_{\varepsilon}$ by:
$K_{\varepsilon}(X,Y,Z):=\frac{1}{2}T_{\varepsilon}(X,Y,Z),\quad X,Y,Z\in\mathfrak{X}(M).$
The metric connection with totally skew-symmetric torsion for parameter $\varepsilon$ is:
$\nabla^{C_{\varepsilon}}_{X}Y:=\nabla^{LC_{\varepsilon}}_{X}Y+K_{\varepsilon}(X,Y,\cdot)^{\sharp_{\varepsilon}},$
where $\nabla^{LC_{\varepsilon}}$ is the Levi-Civita connection of $g_{\varepsilon}$ and $\sharp_{\varepsilon}$ denotes the musical isomorphism $T^{*}M\to TM$ induced by $g_{\varepsilon}$.
Proposition 2.11 (Properties of the connection family).
For each $\varepsilon\in[0,1]$, the connection $\nabla^{C_{\varepsilon}}$ satisfies:
  1. (i)
    $\nabla^{C_{\varepsilon}}g_{\varepsilon}=0$ (metric compatibility).
  2. (ii)
    The torsion tensor of $\nabla^{C_{\varepsilon}}$ is exactly $T_{\varepsilon}$, i.e.:
    $\nabla^{C_{\varepsilon}}_{X}Y-\nabla^{C_{\varepsilon}}_{Y}X-[X,Y]=T_{\varepsilon}(X,Y,\cdot)^{\sharp_{\varepsilon}}.$
  3. (iii)
    For $\varepsilon>0$, $\nabla^{LC_{\varepsilon}}T_{\varepsilon}\neq 0$ (non-parallel torsion), provided $F^{(\pm)}\neq 0$.
Proof.
Properties (i) and (ii) are standard consequences of the definition $\nabla^{C_{\varepsilon}}=\nabla^{LC_{\varepsilon}}+\tfrac{1}{2}T_{\varepsilon}$: adding a totally skew-symmetric $(2,1)$-tensor to the Levi-Civita connection preserves metric compatibility, and the torsion of the resulting connection equals the skew-symmetric tensor (see, e.g., [15]).
For (iii), Appendix A shows that $\nabla^{LC_{\varepsilon}}d\phi_{\pm}=\tfrac{\varepsilon c^{(\pm)}}{2}(e^{2}\otimes e^{1}-e^{1}\otimes e^{2})$, which is nonzero for $\varepsilon>0$ and $c^{(\pm)}\neq 0$. Since $T_{\varepsilon}=c^{(+)}\operatorname{vol}_{\mathrm{BZ}}\wedge(d\phi_{+}+\varepsilon\pi^{*}A^{(+)})+c^{(-)}\operatorname{vol}_{\mathrm{BZ}}\wedge(d\phi_{-}+\varepsilon\pi^{*}A^{(-)})$ and the base form $\operatorname{vol}_{\mathrm{BZ}}$ is parallel, the Leibniz rule gives $\nabla^{LC_{\varepsilon}}T_{\varepsilon}\neq 0$. ∎
2.5 Parallel forms and the obstruction kernel
A crucial ingredient for the PT lower bound is the space of forms that are parallel with respect to the Levi-Civita connection of the product metric ($\varepsilon=0$) and their behavior under deformation.
Definition 2.12 (Spaces of parallel forms on the factors).
For the product metric $g_{0}=g_{\mathrm{BZ}}\oplus d\phi_{+}^{2}\oplus d\phi_{-}^{2}$, we define:
$\mathcal{P}_{2}(T^{2}_{\mathrm{BZ}})$ $:=\{\alpha\in\Omega^{2}(T^{2}_{\mathrm{BZ}})\mid\nabla^{g_{\mathrm{BZ}}}\alpha=0\}=\mathbb{R}\,\operatorname{vol}_{\mathrm{BZ}},$
$\mathcal{P}_{1}(S^{1}_{\phi_{+}}\times S^{1}_{\phi_{-}})$ $:=\{\beta\in\Omega^{1}(S^{1}_{\phi_{+}}\times S^{1}_{\phi_{-}})\mid\nabla^{g_{\text{fibre}}}\beta=0\}=\operatorname{span}\{d\phi_{+},d\phi_{-}\}.$
For $\varepsilon>0$, we define $\mathcal{P}_{1}^{\varepsilon}(S^{1}_{\phi_{+}}\times S^{1}_{\phi_{-}})$ as the space of 1-forms on $M$ that are $g_{\varepsilon}$-parallel and vanish on $\mathcal{H}_{\varepsilon}$ (vertical parallel 1-forms).
Theorem 2.13 (Deformation of vertical parallel 1-forms).
Under Assumption 2.3, for the deformation family $g_{\varepsilon}$, the space of parallel vertical 1-forms $\mathcal{P}_{1}^{\varepsilon}(S^{1}_{\phi_{+}}\times S^{1}_{\phi_{-}})$ satisfies:
  1. 1.
    For $\varepsilon>0$, the space has dimension 1 and consists of constant-coefficient forms $\eta=c_{+}d\phi_{+}+c_{-}d\phi_{-}$ satisfying the algebraic constraint:
[…]
  2. 2.
    For the product limit $\varepsilon=0$, the constraint vanishes and the space has dimension 2, being spanned by $\{d\phi_{+},d\phi_{-}\}$.
Proof.
The condition for a 1-form $\eta$ to be parallel is $\nabla^{LC_{\varepsilon}}\eta=0$. For a vertical form $\eta=c_{+}d\phi_{+}+c_{-}d\phi_{-}$, we evaluate the covariant derivative along horizontal directions. As derived in Appendix A using the connection 1-forms $\omega^{a}_{\;b}$ of the Kaluza–Klein metric $g_{\varepsilon}$, we have:
$\nabla^{LC_{\varepsilon}}\eta=\frac{\varepsilon}{2}(c_{+}c^{(+)}+c_{-}c^{(-)})(e^{2}\otimes e^{1}-e^{1}\otimes e^{2}).$
For any $\varepsilon>0$, the vanishing of this expression requires $c_{+}c^{(+)}+c_{-}c^{(-)}=0$, which is equivalent to $c_{+}+\lambda c_{-}=0$. This constraint defines a 1-dimensional subspace. At the point $\varepsilon=0$, the entire expression vanishes regardless of the coefficients $c_{\pm}$, reflecting the fact that the Levi-Civita connection of the product metric decouples. This completes the proof (see Appendix A for the component-wise tensor derivation). ∎
Remark 2.14 (Global extension of vertical parallel forms).
Since the vertical distribution $\mathcal{V}=\ker d\pi$ is integrable (being the kernel of a submersion) and the fibers $\pi^{-1}(k)$ are totally geodesic flat tori (Theorem 2.7(iv)), any vertical 1-form $\eta$ that is $g_{\varepsilon}$-parallel at one point $p\in M$ extends uniquely by parallel transport to a globally $g_{\varepsilon}$-parallel form. The algebraic constraint $c_{+}+\lambda c_{-}=0$ derived at one point in the proof of Theorem 2.13 therefore holds everywhere, establishing that $\dim\mathcal{P}_{1}^{\varepsilon}(S^{1}_{\phi_{+}}\times S^{1}_{\phi_{-}})=1$ globally for all $\varepsilon>0$.
Definition 2.15 (Obstruction kernel for the product metric).
For the product metric $g_{0}$, let $\mathcal{V}_{2,1}$ be the span of the harmonic representative of the $(2,1)$-Künneth component of $T_{0}]$ (this will be computed explicitly in Section 3). Define:
$\mathcal{K}_{0}:=\left(\mathcal{V}_{2,1}\cap(\mathcal{P}_{2}(T^{2}_{\mathrm{BZ}})\otimes\mathcal{P}_{1}(S^{1}_{\phi_{+}}\times S^{1}_{\phi_{-}}))\right)\oplus\left(\mathcal{V}_{1,2}\cap(\mathcal{P}_{1}(T^{2}_{\mathrm{BZ}})\otimes\mathcal{P}_{2}(S^{1}_{\phi_{+}}\times S^{1}_{\phi_{-}}))\right).$
Lemma 2.16 (Obstruction kernel for $\varepsilon>0$).
For $\varepsilon>0$, let $\mathcal{K}_{\varepsilon}$ be the obstruction kernel defined using the space of parallel vertical 1-forms $\mathcal{P}_{1}^{\varepsilon}(S^{1}_{\phi_{+}}\times S^{1}_{\phi_{-}})$ with respect to the metric $g_{\varepsilon}$. Under Assumption 2.3, we have:
$\mathcal{K}_{\varepsilon}=0\quad\text{for all }\varepsilon>0.$
Proof.
The $(2,1)$-component of the torsion class is represented by a simple tensor that will be explicitly computed in Section 3; it has the form $\xi_{0}=[\operatorname{vol}_{\mathrm{BZ}}]\otimes(a_{+}[d\phi_{+}]+a_{-}[d\phi_{-}])$ with $a_{-}=\lambda a_{+}$. The kernel $\mathcal{K}_{\varepsilon}$ is defined by the intersection of the span of its harmonic representative with the space $\mathcal{P}_{2}(T^{2}_{\mathrm{BZ}})\otimes\mathcal{P}_{1}^{\varepsilon}(S^{1}_{\phi_{+}}\times S^{1}_{\phi_{-}})$.
As established in Theorem 2.13, for $\varepsilon>0$, any parallel vertical 1-form $\eta\in\mathcal{P}_{1}^{\varepsilon}$ must satisfy the algebraic constraint $c_{+}+\lambda c_{-}=0$, where $\lambda=c^{(-)}/c^{(+)}$. However, the harmonic representative has coefficients satisfying $a_{-}=\lambda a_{+}$. Substituting these into the parallel condition yields:
$a_{+}+\lambda a_{-}=a_{+}+\lambda(\lambda a_{+})=a_{+}(1+\lambda^{2}).$
Since $c^{(\pm)}\neq 0$, it follows that $1+\lambda^{2}>0$ (as $\lambda\in\mathbb{R}$). Thus, the parallel condition $a_{+}(1+\lambda^{2})=0$ is satisfied if and only if $a_{+}=0$, which implies $\omega]=0$. For any non-trivial torsion class, the intersection is therefore $\{0\}$, proving that $\mathcal{K}_{\varepsilon}=0$. ∎
Remark 2.17 (Discontinuity of the obstruction kernel).
The dimension of the obstruction kernel jumps at $\varepsilon=0$: we have $\dim\mathcal{K}_{0}=1$ (as will be shown in Section 3) while $\dim\mathcal{K}_{\varepsilon}=0$ for all $\varepsilon>0$. This discontinuity reflects the fact that the geometric product structure at $\varepsilon=0$ allows the mixed class to be completely absorbed by parallel forms, while the Kaluza–Klein coupling for $\varepsilon>0$ imposes additional holonomy constraints that prevent this absorption.
2.6 Local adapted gauge
To perform explicit curvature calculations, we work in local gauges that simplify the metric at a given point.
Definition 2.18 (Adapted gauge at a point).
Let $p\in M$ with $\pi(p)=k_{0}\in T^{2}_{\mathrm{BZ}}$. A gauge is called adapted at $p$ if:
  1. (i)
    $A^{(\pm)}(k_{0})=0$,
  2. (ii)
    $(d\phi_{\pm}+\varepsilon\pi^{*}A^{(\pm)})|_{p}=d\phi_{\pm}|_{p}$,
  3. (iii)
    The metric $g_{\varepsilon}$ satisfies $g_{\varepsilon}|_{p}=dk_{x}^{2}+dk_{y}^{2}+d\phi_{+}^{2}+d\phi_{-}^{2}$.
Lemma 2.19 (Existence of adapted gauge for the family).
For any point $p\in M$ and for every value of the deformation parameter $\varepsilon\in[0,1]$, there exists a local gauge transformation such that the connection $\nabla^{C_{\varepsilon}}$ is adapted at $p$.
Proof.
By Definition 2.5, the connection 1-forms of the family are given by $\Theta^{(\pm)}_{\varepsilon}=d\phi_{\pm}+\varepsilon\pi^{*}A^{(\pm)}$. An adapted gauge at $p$ requires that the connection 1-forms satisfy $\Theta^{(\pm)}_{\varepsilon}|_{p}=d\phi_{\pm}|_{p}$, which is equivalent to the vanishing of the local potentials $\varepsilon A^{(\pm)}$ at the point $p$.
For any $\varepsilon\in[0,1]$, consider the standard gauge transformation $A^{(\pm)}\to A^{(\pm)}+df^{(\pm)}$. By choosing the functions $f^{(\pm)}$ such that $df^{(\pm)}|_{p}=-A^{(\pm)}|_{p}$ (which is always possible in a local neighborhood of $p$), we obtain a transformed potential that vanishes at $p$. Since the parameter $\varepsilon$ scales the potential linearly, the condition $\varepsilon(A^{(\pm)}+df^{(\pm)})|_{p}=0$ is satisfied for all $\varepsilon$. Thus, a single gauge transformation, independent of $\varepsilon$, suffices to make the entire family adapted at $p$. ∎
2.7 Summary of the geometric framework
We have established the following geometric framework:
  1. 1.
    The extended parameter space is the smooth product manifold $M=T^{2}_{\mathrm{BZ}}\times S^{1}_{\phi_{+}}\times S^{1}_{\phi_{-}}$, equipped with a one-parameter family of metrics $\{g_{\varepsilon}\}_{\varepsilon\in[0,1]}$ interpolating between a Riemannian product metric ($\varepsilon=0$) and the physical Kaluza–Klein metric ($\varepsilon=1$).
  2. 2.
    For each $\varepsilon$, we have a metric connection $\nabla^{C_{\varepsilon}}$ with totally skew-symmetric torsion $T_{\varepsilon}$. The cohomology class $T_{\varepsilon}]$ is constant in $\varepsilon$ and admits a Künneth decomposition into mixed components (Lemma 2.9).
  3. 3.
    The space of vertical parallel 1-forms undergoes a discontinuous jump at $\varepsilon=0$: $\dim\mathcal{P}_{1}^{0}=2$ while $\dim\mathcal{P}_{1}^{\varepsilon}=1$ for all $\varepsilon>0$ (Theorem 2.13 and Lemma 2.14).
  4. 4.
    Consequently, the obstruction kernel $\mathcal{K}_{\varepsilon}$ vanishes for all $\varepsilon>0$ (Lemma 2.16), while $\mathcal{K}_{0}$ is non-trivial. This discontinuity is the key to obtaining a non-trivial lower bound for the physical connection.
In the next section, we compute the topological invariants (mixed cohomology class and its tensor rank) explicitly and determine the reduced rank $r^{\sharp}_{\varepsilon}=\operatorname{rank}_{\mathbb{R}}([T_{\varepsilon}]_{\mathrm{mixed}})-\dim\mathcal{K}_{\varepsilon}$ for both $\varepsilon=0$ and $\varepsilon>0$.
3 Cohomology of the Extended Space and the Mixed Tensor Rank
In this section, we analyze the cohomological structure of the extended parameter space $M$ and compute the topological invariants that enter the PT lower bound [12]. Crucially, we distinguish between:
  • Topological invariants (independent of the metric): the cohomology class $\omega]$ and its mixed tensor rank.
  • Geometric data (depending on the metric): the spaces of parallel forms $\mathcal{P}_{k}$ and the obstruction kernel $\mathcal{K}$.
Throughout, we work under Assumption 2.3 (constant Berry curvatures).
3.1 De Rham cohomology and Künneth decomposition
Since $M=T^{2}_{\mathrm{BZ}}\times S^{1}_{\phi_{+}}\times S^{1}_{\phi_{-}}$ is a product manifold (as a smooth manifold), its de Rham cohomology is given by the Künneth theorem:
Theorem 3.1 (Cohomology of $M$).
The de Rham cohomology groups of $M$ are:
$H^{0}(M;\mathbb{R})$ $\cong\mathbb{R},$
$H^{1}(M;\mathbb{R})$ $\cong\mathbb{R}^{4},$
$H^{2}(M;\mathbb{R})$ $\cong\mathbb{R}^{6},$
$H^{3}(M;\mathbb{R})$ $\cong\mathbb{R}^{4},$
$H^{4}(M;\mathbb{R})$ $\cong\mathbb{R}.$
Moreover, the Künneth isomorphism gives a canonical decomposition:
$H^{3}(M;\mathbb{R})\cong\bigoplus_{p+q=3}H^{p}(T^{2}_{\mathrm{BZ}};\mathbb{R})\otimes H^{q}(S^{1}_{\phi_{+}}\times S^{1}_{\phi_{-}};\mathbb{R}).$
Explicitly, the non-zero summands are:
$H^{3}(M;\mathbb{R})\cong\left(H^{2}(T^{2}_{\mathrm{BZ}})\otimes H^{1}(S^{1}_{\phi_{+}}\times S^{1}_{\phi_{-}})\right)\oplus\left(H^{1}(T^{2}_{\mathrm{BZ}})\otimes H^{2}(S^{1}_{\phi_{+}}\times S^{1}_{\phi_{-}})\right).$
The torsion 3-form $T$ defined in Definition 2.8 (for $\varepsilon=1$) represents a well-defined cohomology class:
Definition 3.2 (Torsion cohomology class).
The torsion cohomology class is:
$\omega]:=[T]\in H^{3}(M;\mathbb{R}).$
Under the deformation family $g_{\varepsilon}$, the class $\omega]$ is constant in $\varepsilon$, as shown in Lemma 2.9.
3.2 Harmonic representative for the product metric
For the product metric $g_{0}=g_{\mathrm{BZ}}\oplus d\phi_{+}^{2}\oplus d\phi_{-}^{2}$, the harmonic representative of $\omega]$ can be written explicitly.
Theorem 3.3 (Harmonic representative for $g_{0}$).
Under Assumption 2.3, the unique $g_{0}$-harmonic representative of $\omega]$ is:
$\omega_{h}^{0}=\operatorname{vol}_{\mathrm{BZ}}\wedge(a_{+}d\phi_{+}+a_{-}d\phi_{-}),$
where the constants $a_{\pm}\in\mathbb{R}$ satisfy:
$a_{-}=\lambda a_{+},\quad\lambda=\frac{c^{(-)}}{c^{(+)}}.$
These constants are uniquely determined by the cohomology classes $F^{(\pm)}]$.
Proof.
Since $g_{0}$ is a product metric, the Hodge Laplacian decouples. The form $\eta:=\operatorname{vol}_{\mathrm{BZ}}\wedge(a_{+}d\phi_{+}+a_{-}d\phi_{-})$ is closed because $\operatorname{vol}_{\mathrm{BZ}}$ is closed and $d\phi_{\pm}$ are closed.
It is coclosed if and only if $\delta\eta=0$. In the product metric, using $\eta=a_{+}(*_{\mathrm{BZ}}\operatorname{vol}_{\mathrm{BZ}})\wedge(*_{\text{fibre}}d\phi_{+})+a_{-}(*_{\mathrm{BZ}}\operatorname{vol}_{\mathrm{BZ}})\wedge(*_{\text{fibre}}d\phi_{-})$, where $_{\mathrm{BZ}}\operatorname{vol}_{\mathrm{BZ}}=1$ and $_{\text{fibre}}d\phi_{\pm}=\pm d\phi_{\mp}$, the coclosed condition gives:
$a_{+}c^{(+)}+a_{-}c^{(-)}=0,$
which is equivalent to $a_{-}=\lambda a_{+}$. Uniqueness follows from the Hodge theorem. ∎
Remark 3.4 (Harmonic representative for $\varepsilon>0$).
For $\varepsilon>0$, the harmonic representative $\omega_{h}^{\varepsilon}$ with respect to $g_{\varepsilon}$ is not a wedge product of forms on the factors, because the metric is not a product. However, its cohomology class is the same as $\omega_{h}^{0}$.
3.3 Mixed tensor rank
We now define the mixed tensor rank for a cohomology class in $H^{3}(M)$. This is a purely topological invariant.
Definition 3.5 (Mixed tensor rank).
Let $\eta]\in H^{3}(M;\mathbb{R})$. Under the Künneth decomposition, write
$\eta]=\sum_{i=1}^{r_{2,1}}[\alpha_{i}]\otimes[\beta_{i}]+\sum_{j=1}^{r_{1,2}}[\tilde{\alpha}_{j}]\otimes[\tilde{\beta}_{j}],$
where $\alpha_{i}]\in H^{2}(T^{2}_{\mathrm{BZ}})$, $\beta_{i}]\in H^{1}(S^{1}_{\phi_{+}}\times S^{1}_{\phi_{-}})$, and similarly for the $(1,2)$-part. The mixed tensor rank of $\eta]$ is
$\operatorname{rank}_{\mathbb{R}}([\eta]_{\mathrm{mixed}}):=\min\{r_{2,1}+r_{1,2}\},$
where the minimum is taken over all such representations.
Theorem 3.6 (Mixed tensor rank of the torsion class).
Under Assumption 2.3, the mixed tensor rank of $\omega]$ is
$\operatorname{rank}_{\mathbb{R}}([\omega]_{\mathrm{mixed}})=1.$
Proof.
From the explicit form of the harmonic representative $\omega_{h}^{0}$ (Theorem 3.3), we have
$\omega]=[\operatorname{vol}_{\mathrm{BZ}}]\otimes\left(a_{+}[d\phi_{+}]+a_{-}[d\phi_{-}]\right).$
Since $a_{+}[d\phi_{+}]+a_{-}[d\phi_{-}]$ is a single element in $H^{1}(S^{1}_{\phi_{+}}\times S^{1}_{\phi_{-}})$, this shows that the mixed rank is at most 1. It cannot be 0 because $\omega]\neq 0$ (under Assumption 2.3, $c^{(\pm)}\neq 0$). ∎
Remark 3.7 (Geometric interpretation).
The rank being 1 indicates that the topological coupling between the Brillouin zone and the phase directions is irreducible: the torsion class cannot be decomposed into two independent tensor products. This reflects the fact that the two U(1) sectors are not independent topologically when their curvatures are proportional.
3.4 Parallel forms and obstruction kernel for the product metric
To apply the PT theorem, we need the spaces of parallel forms for the product metric $g_{0}$.
Lemma 3.8 (Parallel forms on the factors for $g_{0}$).
For the product metric $g_{0}=g_{\mathrm{BZ}}\oplus d\phi_{+}^{2}\oplus d\phi_{-}^{2}$, we have:
$\mathcal{P}_{2}(T^{2}_{\mathrm{BZ}},g_{\mathrm{BZ}})$ $=\mathbb{R}\,\operatorname{vol}_{\mathrm{BZ}},$
$\mathcal{P}_{1}(S^{1}_{\phi_{+}}\times S^{1}_{\phi_{-}},g_{0}|_{\text{fibre}})$ $=\operatorname{span}\{d\phi_{+},d\phi_{-}\}.$
Proof.
The base is flat, so all harmonic forms are parallel. The fibre is a flat torus, so all constant 1-forms are parallel. ∎
Now we compute the obstruction kernel $\mathcal{K}_{0}$ for the product metric.
Definition 3.9 (Obstruction kernel for $g_{0}$).
Let $\mathcal{V}_{2,1}$ be the subspace of $H^{2}(T^{2}_{\mathrm{BZ}})\otimes H^{1}(S^{1}_{\phi_{+}}\times S^{1}_{\phi_{-}})$ spanned by the harmonic representatives of the $(2,1)$-component of $\omega]$ with respect to $g_{0}$ (as given in Definition 3.5). Define
$\mathcal{K}_{0}:=\left(\mathcal{V}_{2,1}\cap(\mathcal{P}_{2}(T^{2}_{\mathrm{BZ}})\otimes\mathcal{P}_{1}(S^{1}_{\phi_{+}}\times S^{1}_{\phi_{-}}))\right)\oplus\left(\mathcal{V}_{1,2}\cap(\mathcal{P}_{1}(T^{2}_{\mathrm{BZ}})\otimes\mathcal{P}_{2}(S^{1}_{\phi_{+}}\times S^{1}_{\phi_{-}}))\right).$
Theorem 3.10 (Obstruction kernel for $g_{0}$).
Under Assumption 2.3, for the product metric $g_{0}$, we have
$\dim\mathcal{K}_{0}=1.$
Proof.
From Theorem 3.3, the harmonic representative of the $(2,1)$-component is
$\xi_{0}=\operatorname{vol}_{\mathrm{BZ}}\otimes(a_{+}d\phi_{+}+a_{-}d\phi_{-}).$
Thus $\mathcal{V}_{2,1}=\operatorname{span}\{\xi_{0}\}$. Since $\operatorname{vol}_{\mathrm{BZ}}\in\mathcal{P}_{2}(T^{2}_{\mathrm{BZ}})$ and $a_{+}d\phi_{+}+a_{-}d\phi_{-}\in\mathcal{P}_{1}(S^{1}_{\phi_{+}}\times S^{1}_{\phi_{-}})$ (it is a constant linear combination), we have
$\xi_{0}\in\mathcal{P}_{2}(T^{2}_{\mathrm{BZ}})\otimes\mathcal{P}_{1}(S^{1}_{\phi_{+}}\times S^{1}_{\phi_{-}}).$
Hence
$\mathcal{V}_{2,1}\subseteq\mathcal{P}_{2}(T^{2}_{\mathrm{BZ}})\otimes\mathcal{P}_{1}(S^{1}_{\phi_{+}}\times S^{1}_{\phi_{-}}),$
so the intersection is all of $\mathcal{V}_{2,1}$, and $\dim\mathcal{K}_{0}=\dim\mathcal{V}_{2,1}=1$. The $(1,2)$-component is zero (since $b_{1}(T^{2}_{\mathrm{BZ}})\cdot b_{2}(S^{1}\times S^{1})=0$), so it contributes nothing. ∎
3.5 Reduced rank for the product metric
We now compute the key quantity that appears in the PT lower bound for the product metric.
Definition 3.11 (Reduced rank for $g_{0}$).
The reduced rank for the product metric $g_{0}$ is
$r_{0}^{\sharp}:=\operatorname{rank}_{\mathbb{R}}([\omega]_{\mathrm{mixed}})-\dim\mathcal{K}_{0}.$
Corollary 3.12 (Value of $r_{0}^{\sharp}$).
Under Assumption 2.3,
$r_{0}^{\sharp}=0.$
Proof.
By Theorem 3.6, $\operatorname{rank}_{\mathbb{R}}([\omega]_{\mathrm{mixed}})=1$. By Theorem 3.10, $\dim\mathcal{K}_{0}=1$. Hence $r_{0}^{\sharp}=1-1=0$. ∎
Remark 3.13 (Interpretation).
For the product metric $g_{0}$, the PT theorem gives the lower bound
$\dim\mathfrak{hol}^{\mathrm{off}}(\nabla^{C_{0}})\geq r_{0}^{\sharp}=0,$
which is trivial. This is expected because for $\varepsilon=0$, the connection $\nabla^{C_{0}}$ is adapted to the product structure and its curvature may be block-diagonal. The non-trivial bound will arise from the deformation to $\varepsilon>0$.
3.6 Behavior under deformation
We now examine how the obstruction kernel changes as we move away from the product metric.
Theorem 3.14 (Obstruction kernel for $\varepsilon>0$).
For $\varepsilon>0$, under Assumption 2.3, the obstruction kernel $\mathcal{K}_{\varepsilon}$ (defined analogously using the spaces of parallel forms for $g_{\varepsilon}$) satisfies
$\mathcal{K}_{\varepsilon}=0.$
Consequently, $\dim\mathcal{K}_{\varepsilon}=0$.
Proof.
The $(2,1)$-component of the torsion class is represented by the simple tensor $\xi_{0}=[\operatorname{vol}_{\mathrm{BZ}}]\otimes(a_{+}[d\phi_{+}]+a_{-}[d\phi_{-}])$ (Theorem 3.3). The kernel $\mathcal{K}_{\varepsilon}$ is defined by the intersection of the span of its harmonic representative with the space $\mathcal{P}_{2}(T^{2}_{\mathrm{BZ}})\otimes\mathcal{P}_{1}^{\varepsilon}(S^{1}_{\phi_{+}}\times S^{1}_{\phi_{-}})$.
As established in Theorem 2.13, for $\varepsilon>0$, any parallel vertical 1-form $\eta\in\mathcal{P}_{1}^{\varepsilon}$ must satisfy the algebraic constraint $c_{+}+\lambda c_{-}=0$, where $\lambda=c^{(-)}/c^{(+)}$. However, the harmonic representative of $\omega]$ has coefficients satisfying $a_{-}=\lambda a_{+}$. Substituting these into the parallel condition yields:
$a_{+}+\lambda a_{-}=a_{+}+\lambda(\lambda a_{+})=a_{+}(1+\lambda^{2}).$
Since $c^{(\pm)}\neq 0$ (Assumption 2.3), it follows that $1+\lambda^{2}>0$ for any real $\lambda$. Thus, the parallel condition $a_{+}(1+\lambda^{2})=0$ is satisfied if and only if $a_{+}=0$, which implies $\omega]=0$. For any non-trivial torsion class, the intersection is therefore $\{0\}$, proving that $\mathcal{K}_{\varepsilon}=0$. ∎
Corollary 3.15 (Reduced rank for $\varepsilon>0$).
For $\varepsilon>0$, the reduced rank $r_{\varepsilon}^{\sharp}:=\operatorname{rank}_{\mathbb{R}}([\omega]_{\mathrm{mixed}})-\dim\mathcal{K}_{\varepsilon}$ satisfies
$r_{\varepsilon}^{\sharp}=1.$
Proof.
By Theorem 3.6, $\operatorname{rank}_{\mathbb{R}}([\omega]_{\mathrm{mixed}})=1$. By Theorem 3.14, $\dim\mathcal{K}_{\varepsilon}=0$ for $\varepsilon>0$. Hence $r_{\varepsilon}^{\sharp}=1-0=1$. ∎
3.7 Summary
We have established the following topological and geometric data:
  1. 1.
    The torsion class $\omega]\in H^{3}(M;\mathbb{R})$ has mixed tensor rank $\operatorname{rank}_{\mathbb{R}}([\omega]_{\mathrm{mixed}})=1$ (Theorem 3.6), a topological invariant.
  2. 2.
    For the product metric $g_{0}$, the obstruction kernel $\mathcal{K}_{0}$ has dimension 1, giving reduced rank $r_{0}^{\sharp}=0$ (Corollary 3.12).
  3. 3.
    For the deformed metrics $g_{\varepsilon}$ with $\varepsilon>0$, the obstruction kernel vanishes, giving reduced rank $r_{\varepsilon}^{\sharp}=1$ (Corollary 3.15).
  4. 4.
    The PT theorem applies directly to $(M,g_{0},\nabla^{C_{0}})$, yielding
    $\dim\mathfrak{hol}^{\mathrm{off}}(\nabla^{C_{0}})\geq r_{0}^{\sharp}=0.$
    This bound is trivial, but for $\varepsilon>0$ the bound becomes non-trivial, as will be established in Section 4 through direct curvature analysis.
In the next section, we apply the deformation strategy and establish the non-trivial lower bound for the physical connection $\nabla^{C}=\nabla^{C_{1}}$.
4 Applying the PT Theorem via Deformation and Direct Analysis
The PT lower bound [12] applies directly to genuine Riemannian product manifolds. Our physical metric $g_{M}$ is a Kaluza–Klein metric that is not a product. However, we have constructed a smooth deformation $g_{\varepsilon}$ (Definition 2.5) that interpolates between the product metric $g_{0}$ and the physical metric $g_{M}=g_{1}$. In this section, we:
  1. 1.
    Apply the original PT theorem to the product metric $g_{0}$ at $\varepsilon=0$.
  2. 2.
    Establish the lower bound for all $\varepsilon>0$ through direct curvature analysis in adapted gauge, working with the torsion form explicitly.
  3. 3.
    Apply the result to the physical connection at $\varepsilon=1$.
The key observation is that while the PT bound is trivial at the product limit ($\varepsilon=0$), it becomes non-trivial for all $\varepsilon>0$ due to the vanishing of the obstruction kernel (Lemma 2.16). This allows us to establish the desired inequality for the physical system through explicit parametric calculation in the constants $c^{(\pm)}$ characterizing the SOC BEC model.
4.1 Applying the PT theorem to the product metric $g_{0}$
For $\varepsilon=0$, the metric $g_{0}=g_{\mathrm{BZ}}\oplus d\phi_{+}^{2}\oplus d\phi_{-}^{2}$ is a Riemannian product. The corresponding connection $\nabla^{C_{0}}$ has torsion $T_{0}=F^{(+)}\wedge d\phi_{+}+F^{(-)}\wedge d\phi_{-}$. The PT theorem can be applied directly to this geometric setup.
Theorem 4.1 (PT bound for the product metric).
For the product metric $g_{0}$ and the connection $\nabla^{C_{0}}$, the PT theorem gives
$\dim\mathfrak{hol}^{\mathrm{off}}(\nabla^{C_{0}})\;\geq\;r^{\sharp}_{0}\;=\;0.$
Proof.
The PT theorem (Theorem 5.2 of [12]) states that
$\dim\mathfrak{hol}^{\mathrm{off}}(\nabla^{C_{0}})\;\geq\;\operatorname{rank}_{\mathbb{R}}([\omega]_{\mathrm{mixed}})-\dim\mathcal{K}_{0}.$
From Corollary 3.12 (Section 3) we have
$r^{\sharp}_{0}=\min(\operatorname{rank}_{\mathbb{R}}([\omega])-\dim\mathcal{K}_{0},1)=\min(1-1,1)=0.$
Remark 4.2 (Triviality at the product limit).
The bound is trivial for the product metric, as expected. In the product geometry, the connection $\nabla^{C_{0}}$ may have block-diagonal curvature because the topological obstruction encoded by the mixed class is completely absorbed by the parallel forms: the harmonic representative of $\omega]$ belongs to $\mathcal{P}_{2}(T^{2}_{\mathrm{BZ}})\otimes\mathcal{P}_{1}(S^{1}_{\phi_{+}}\times S^{1}_{\phi_{-}})$, yielding $\dim\mathcal{K}_{0}=1$ and hence $r^{\sharp}_{0}=0$. This absorption is the geometric reason for the vanishing of the lower bound at $\varepsilon=0$.
4.2 Direct curvature analysis for $\varepsilon>0$
For $\varepsilon>0$, the metric $g_{\varepsilon}$ is no longer a Riemannian product, and the PT theorem does not apply directly. However, we can establish the lower bound through explicit pointwise computation of the off-diagonal curvature components for the SOC BEC model.
The key geometric fact is that for $\varepsilon>0$, the space of vertical parallel 1-forms drops from dimension 2 to dimension 1 (Theorem 2.13), causing the obstruction kernel to vanish: $\mathcal{K}_{\varepsilon}=0$ (Lemma 2.16). This yields a non-trivial reduced rank $r^{\sharp}_{\varepsilon}=1$ for all $\varepsilon>0$.
Definition 4.3 (Off-diagonal curvature test for $\varepsilon>0$).
Fix $\varepsilon>0$ and a point $p\in M$. Let $\mathcal{H}_{p}$ and $\mathcal{V}_{p}$ denote the horizontal and vertical subspaces of $T_{p}M$, respectively, and let
$\pi_{\mathrm{off}}:\mathfrak{so}(T_{p}M)\;\longrightarrow\;\mathrm{Hom}(\mathcal{H}_{p},\mathcal{V}_{p})\oplus\mathrm{Hom}(\mathcal{V}_{p},\mathcal{H}_{p})$
be the projection onto the off-diagonal component (endomorphisms that map $\mathcal{H}_{p}\to\mathcal{V}_{p}$ and $\mathcal{V}_{p}\to\mathcal{H}_{p}$).
Define the off-diagonal curvature test at $p$ as the collection of endomorphisms
$\mathcal{R}^{\mathrm{off}}_{p}(\varepsilon)\;:=\;\bigl\{\,\pi_{\mathrm{off}}\!\bigl(R^{C_{\varepsilon}}(X,Z)\bigr)\;\big|\;X\in\mathcal{H}_{p},\;Z\in\mathcal{V}_{p}\,\bigr\}\;\subset\;\mathrm{Hom}(\mathcal{H}_{p},\mathcal{V}_{p})\oplus\mathrm{Hom}(\mathcal{V}_{p},\mathcal{H}_{p}).$
The off-diagonal curvature is non-trivial at $p$ if $\mathcal{R}^{\mathrm{off}}_{p}(\varepsilon)\neq\{0\}$, i.e. if at least one mixed-input curvature operator has a nonzero off-diagonal projection.
Remark 4.4 (Intrinsic nature of the off-diagonal curvature test).
Although explicit computations are performed in an adapted gauge (Definition 2.18), the set $\mathcal{R}^{\mathrm{off}}_{p}(\varepsilon)$ is intrinsically defined because:
  1. 1.
    The curvature tensor $R^{C_{\varepsilon}}(X,Z)$ is a $(1,3)$-tensor, independent of any choice of coordinates or gauge.
  2. 2.
    The splitting $T_{p}M=\mathcal{H}_{p}\oplus\mathcal{V}_{p}$ is determined by the Kaluza–Klein structure: $\mathcal{V}_{p}=\ker d\pi|_{p}$ and $\mathcal{H}_{p}=\mathcal{V}_{p}^{\perp_{g_{\varepsilon}}}$.
  3. 3.
    The projection $\pi_{\mathrm{off}}$ depends only on the splitting, not on the choice of basis within each factor.
By the Ambrose–Singer theorem, if $\mathcal{R}^{\mathrm{off}}_{p}(\varepsilon)\neq\{0\}$, then $\mathfrak{hol}^{\mathrm{off}}_{p}(\nabla^{C_{\varepsilon}})\neq\{0\}$ and hence $\dim\mathfrak{hol}^{\mathrm{off}}_{p}(\nabla^{C_{\varepsilon}})\geq 1$.
Theorem 4.5 (Off-diagonal holonomy bound for SOC BECs).
Consider a two-component SOC BEC with extended parameter space $M=T^{2}_{\mathrm{BZ}}\times S^{1}_{\phi_{+}}\times S^{1}_{\phi_{-}}$ equipped with the Kaluza–Klein metric $g_{\varepsilon}$ for $\varepsilon>0$.
Under Assumption 2.3 (constant Berry curvatures $F^{(\pm)}=c^{(\pm)}\operatorname{vol}_{\mathrm{BZ}}$ with $c^{(\pm)}\neq 0$), the off-diagonal holonomy dimension satisfies:
$\dim\mathfrak{hol}^{\mathrm{off}}(\nabla^{C_{\varepsilon}})\geq 1\quad\text{at every point }p\in M.$ (4.1)
Proof.
We establish the bound by exhibiting a nonzero off-diagonal curvature operator at every point, using the curvature evaluated on mixed inputs (one horizontal, one vertical).
1. Setup in adapted gauge. Fix an arbitrary point $p\in M$ and choose adapted gauge where $A^{(\pm)}|_{p}=0$ (Lemma 2.19). At $p$, the orthonormal frame is
$e_{1}=\partial_{k_{x}},\quad e_{2}=\partial_{k_{y}}\quad(\text{horizontal, }\mathcal{H}_{p}),\qquad e_{3}=\partial_{\phi_{+}},\quad e_{4}=\partial_{\phi_{-}}\quad(\text{vertical, }\mathcal{V}_{p}),$
and the metric reduces to $g_{\varepsilon}|_{p}=dk_{x}^{2}+dk_{y}^{2}+d\phi_{+}^{2}+d\phi_{-}^{2}$.
2. Curvature decomposition for mixed inputs. The curvature of $\nabla^{C_{\varepsilon}}=\nabla^{\mathrm{LC}(\varepsilon)}+\tfrac{1}{2}T_{\varepsilon}$ decomposes as
$R^{C_{\varepsilon}}(X,Y)\;=\;R^{\mathrm{LC}(\varepsilon)}(X,Y)\;+\;\tfrac{1}{2}\bigl[(\nabla^{\mathrm{LC}(\varepsilon)}_{X}T_{\varepsilon})(Y,\cdot)-(\nabla^{\mathrm{LC}(\varepsilon)}_{Y}T_{\varepsilon})(X,\cdot)\bigr]^{\sharp}\;+\;\tfrac{1}{4}[T_{X},T_{Y}],$ (4.2)
where $T_{X}(Z):=T_{\varepsilon}(X,Z,\cdot)^{\sharp}$.
To detect off-diagonal holonomy, we evaluate (4.2) on the mixed pair $X=e_{1}\in\mathcal{H}_{p}$, $Y=e_{3}\in\mathcal{V}_{p}$ and extract the off-diagonal projection
$\pi_{\mathrm{off}}\!\bigl(R^{C_{\varepsilon}}(e_{1},e_{3})\bigr)\;\in\;\mathrm{Hom}(\mathcal{H}_{p},\mathcal{V}_{p})\oplus\mathrm{Hom}(\mathcal{V}_{p},\mathcal{H}_{p}),$
which maps horizontal directions to vertical ones and vice versa.
3. Torsion endomorphisms at $p$. At $p$ in adapted gauge, the torsion is $T_{\varepsilon}|_{p}=c^{(+)}e^{1}\!\wedge e^{2}\!\wedge e^{3}+c^{(-)}e^{1}\!\wedge e^{2}\!\wedge e^{4}$, which is independent of $\varepsilon$. The torsion endomorphism table (Lemma B.4 of Appendix B) gives, in particular:
$T_{e_{1}}(e_{2})$ $=c^{(+)}e_{3}+c^{(-)}e_{4}$ $(\mathcal{H}\to\mathcal{V}),$
$T_{e_{1}}(e_{3})$ $=-c^{(+)}e_{2}$ $(\mathcal{V}\to\mathcal{H}),$
$T_{e_{3}}(e_{1})$ $=c^{(+)}e_{2},\qquad T_{e_{3}}(e_{2})=-c^{(+)}e_{1}$ $(\mathcal{H}\to\mathcal{H}).$
4. Leading contribution: the quadratic torsion term. The commutator $T_{e_{1}},T_{e_{3}}]$ is computed entry-by-entry in Proposition B.6 of Appendix B:
$[T_{e_{1}},T_{e_{3}}](e_{1})$ $=(c^{(+)})^{2}\,e_{3}+c^{(+)}c^{(-)}\,e_{4}$ $(\mathcal{H}\to\mathcal{V}),$ (4.3)
$[T_{e_{1}},T_{e_{3}}](e_{3})$ $=-(c^{(+)})^{2}\,e_{1}$ $(\mathcal{V}\to\mathcal{H}),$ (4.4)
$[T_{e_{1}},T_{e_{3}}](e_{4})$ $=-c^{(+)}c^{(-)}\,e_{1}$ $(\mathcal{V}\to\mathcal{H}).$ (4.5)
Every nonzero entry is off-diagonal: it maps $\mathcal{H}\to\mathcal{V}$ or $\mathcal{V}\to\mathcal{H}$. In particular, (4.4) gives
$\pi_{\mathrm{off}}\!\bigl(\tfrac{1}{4}[T_{e_{1}},T_{e_{3}}]\bigr)(e_{3})=-\frac{(c^{(+)})^{2}}{4}\,e_{1}\;\neq\;0\quad\text{for all }c^{(+)}\neq 0.$ (4.6)
5. Subdominance of the remaining terms. The other two contributions to $R^{C_{\varepsilon}}(e_{1},e_{3})$ are:
  • Levi-Civita curvature: All curvature 2-forms $\Omega^{a}{}_{b}$ of $g_{\varepsilon}$ are proportional to $e^{1}\!\wedge e^{2}$ (Appendix A). Evaluating on the mixed pair $(e_{1},e_{3})$ gives zero: $R^{\mathrm{LC}(\varepsilon)}(e_{1},e_{3})=0$.
  • Torsion derivative: Since the torsion has constant coefficients in adapted gauge and the connection coefficients are $\mathcal{O}(\varepsilon)$, the covariant derivative $\nabla^{\mathrm{LC}(\varepsilon)}T_{\varepsilon}$ is $\mathcal{O}(\varepsilon)$.
Therefore:
$\pi_{\mathrm{off}}\!\bigl(R^{C_{\varepsilon}}(e_{1},e_{3})\bigr)\;=\;\pi_{\mathrm{off}}\!\bigl(\tfrac{1}{4}[T_{e_{1}},T_{e_{3}}]\bigr)\;+\;\mathcal{O}(\varepsilon).$ (4.7)
The leading term (4.6) is a nonzero constant independent of $\varepsilon$, while the correction is continuous and vanishes at $\varepsilon=0$. Hence the total is nonzero for all $\varepsilon\in[0,1]$.
6. Holonomy generation. By the Ambrose–Singer theorem, the holonomy algebra $\mathfrak{hol}_{p}(\nabla^{C_{\varepsilon}})$ contains all curvature endomorphisms $R^{C_{\varepsilon}}(X,Y)$ for $X,Y\in T_{p}M$. Since $\pi_{\mathrm{off}}(R^{C_{\varepsilon}}(e_{1},e_{3}))\neq 0$, there exists at least one nonzero element in the off-diagonal component
$\mathfrak{hol}^{\mathrm{off}}_{p}(\nabla^{C_{\varepsilon}}):=\mathfrak{hol}_{p}(\nabla^{C_{\varepsilon}})\cap\bigl(\mathrm{Hom}(\mathcal{H}_{p},\mathcal{V}_{p})\oplus\mathrm{Hom}(\mathcal{V}_{p},\mathcal{H}_{p})\bigr).$
By homogeneity of the model (Proposition B.2, Appendix B), the curvature components are independent of $p$, so the bound holds globally:
$\dim\mathfrak{hol}^{\mathrm{off}}_{p}(\nabla^{C_{\varepsilon}})\;\geq\;1\qquad\text{for every }p\in M,\ \varepsilon>0,\ c^{(\pm)}\neq 0.$ (4.8)
Corollary 4.6 (Lower bound for $\varepsilon>0$).
For every $\varepsilon>0$ and every point $p\in M$,
$\dim\mathfrak{hol}^{\mathrm{off}}(\nabla^{C_{\varepsilon}})\;\geq\;1.$
Proof.
This is the direct statement of Theorem 4.5. ∎
4.3 Application to the physical connection
Theorem 4.7 (Lower bound for the physical connection).
Under Assumption 2.3, the physical connection $\nabla^{C}=\nabla^{C_{1}}$ satisfies
$\dim\mathfrak{hol}^{\mathrm{off}}(\nabla^{C})\;\geq\;1\qquad\text{for all }p\in M.$
Proof.
This follows immediately from Theorem 4.5 by setting $\varepsilon=1$. The physical connection $\nabla^{C}$ is precisely $\nabla^{C_{1}}$, which satisfies the inequality for every point in $M$. ∎
Remark 4.8 (Direct nature of the argument).
The proof of Theorem 4.7 establishes the bound through direct pointwise computation of the curvature as an explicit function of the SOC BEC model parameters $(c^{(+)},c^{(-)})$. The vanishing of the obstruction kernel for $\varepsilon>0$ (Lemma 2.16) provides the crucial geometric input that makes the bound non-trivial, contrasting sharply with the trivial bound at $\varepsilon=0$.
4.4 Geometric and physical interpretation
The inequality $\dim\mathfrak{hol}^{\mathrm{off}}(\nabla^{C})\geq 1$ has a direct geometric meaning: the curvature of the synthetic gauge connection cannot be made block-diagonal with respect to the splitting $T_{p}M=\mathcal{H}_{p}\oplus\mathcal{V}_{p}$. At least one independent off-diagonal curvature operator persists, mixing momentum directions (crystal momenta $k_{x},k_{y}$) with the phase directions ($\phi_{+},\phi_{-}$). This represents a topological obstruction to flattening the Berry curvature simultaneously along both phase circles.
Remark 4.9 (Topological locking of degrees of freedom).
The off-diagonal holonomy can be interpreted as a form of topological locking: the momentum and phase degrees of freedom are entangled by the mixed cohomology class $\omega]$ in a way that cannot be undone by smooth gauge transformations or metric deformations preserving the cohomology classes $F^{(\pm)}]$. This locking persists even when local gauge transformations can eliminate the Berry curvature in restricted subspaces of the parameter space.
Remark 4.10 (Independence of the total Chern number).
The bound remains non-trivial even when the total Chern number vanishes:
$c_{1}^{\mathrm{tot}}:=c_{1}^{(+)}+c_{1}^{(-)}=\frac{1}{2\pi}\int_{T^{2}_{\mathrm{BZ}}}\left(F^{(+)}+F^{(-)}\right)=0.$
As long as the individual curvatures are non-zero ($c^{(\pm)}\neq 0$), the mixed rank remains $r=1$ and the obstruction kernel vanishes for $\varepsilon>0$, yielding $r^{\sharp}=1$. This demonstrates that the mixed-tensor-rank invariant captures a local topological constraint that persists when global topological charges cancel. The structure of the mixed class—not just its integral—determines the irreducibility of the holonomy.
Remark 4.11 (Robustness under perturbations).
The bound depends only on:
  1. 1.
    The cohomology class $\omega]=[F^{(+)}]\otimes[d\phi_{+}]+[F^{(-)}]\otimes[d\phi_{-}]$, which is a topological invariant.
  2. 2.
    The condition that $c^{(\pm)}\neq 0$, which ensures non-vanishing curvature.
Both properties are stable under small perturbations of the system parameters (Raman laser intensities, trapping frequencies, interaction strengths). Hence, the topological obstruction certified by $\dim\mathfrak{hol}^{\mathrm{off}}\geq 1$ is robust against experimental imperfections and represents a generic feature of the SOC BEC phase diagram.
Remark 4.12 (Relation to the cohomological lower bound of [12]).
The lower bound of [12] applies to cohomologically calibrated connections on Riemannian product manifolds and yields the invariant $r^{\sharp}=\min(r-\dim\mathcal{K},\,1)$, where $r$ is the mixed tensor rank and $\mathcal{K}$ is the obstruction kernel determined by parallel-form strata. In the product limit $\varepsilon=0$ of our model, the metric $g_{0}$ is a Riemannian product on $T^{2}_{\mathrm{BZ}}\times T^{2}_{\mathrm{fiber}}$ with both factors flat, so $\mathcal{P}_{1}(T^{2}_{\mathrm{fiber}})=\mathbb{R}^{2}$ and $r^{\sharp}=0$: the bound is trivial, consistently with the fact that the product connection has reducible holonomy.
For $\varepsilon>0$, the Kaluza–Klein metric $g_{\varepsilon}$ is no longer a Riemannian product, and the theorem of [12] does not apply directly. Nevertheless, the algebraic mechanism that drives the transition $r^{\sharp}=0\to r^{\sharp}=1$ in [12]—namely, the collapse of the parallel-form stratum $\mathcal{P}_{1}$ from $\mathbb{R}^{2}$ to a proper subspace—has an exact counterpart in our setting: Proposition LABEL:prop:parallel-drop shows that $\dim\mathcal{P}_{1}^{\varepsilon}$ drops from $2$ to $1$ as soon as $\varepsilon>0$. The direct curvature computation of Theorem 4.5 then confirms that this drop is indeed accompanied by the appearance of off-diagonal holonomy, as the cohomological framework would predict.
This suggests that the cohomological invariant $r^{\sharp}$ of [12] may capture the relevant obstruction beyond the Riemannian product setting, at least for Kaluza–Klein deformations of product metrics where the parallel-form strata remain computable. A rigorous extension of the lower bound to non-product metrics is an open problem.
4.5 Summary of the proof strategy
The logical structure of the proof can be summarized as follows:
  1. 1.
    Deformation family (Section 2): We construct a smooth one-parameter family of metrics $\{g_{\varepsilon}\}_{\varepsilon\in[0,1]}$ interpolating between the Riemannian product metric $g_{0}=g_{\mathrm{BZ}}\oplus d\phi_{+}^{2}\oplus d\phi_{-}^{2}$ and the physical Kaluza–Klein metric $g_{M}=g_{1}$.
  2. 2.
    Topological invariants (Section 3): We compute the mixed tensor rank $r=\operatorname{rank}_{\mathbb{R}}([\omega]_{\mathrm{mixed}})=1$ and the obstruction kernels:
    $\dim\mathcal{K}_{0}=1\quad(\varepsilon=0),\qquad\dim\mathcal{K}_{\varepsilon}=0\quad(\varepsilon>0).$
    This yields reduced ranks $r^{\sharp}_{0}=0$ (trivial) and $r^{\sharp}_{\varepsilon}=1$ (non-trivial).
  3. 3.
    Product case ($\varepsilon=0$): The PT theorem applies directly, giving the trivial bound $\dim\mathfrak{hol}^{\mathrm{off}}(\nabla^{C_{0}})\geq 0$.
  4. 4.
    Deformed cases ($\varepsilon>0$): Through explicit computation of the curvature $R^{C_{\varepsilon}}(X,Z)$ for mixed inputs $X\in\mathcal{H}_{p}$, $Z\in\mathcal{V}_{p}$ (Theorem 4.5, Appendix B), we show that the off-diagonal projection $\pi_{\mathrm{off}}(R^{C_{\varepsilon}}(X,Z))\neq 0$ at every point, establishing $\dim\mathfrak{hol}^{\mathrm{off}}(\nabla^{C_{\varepsilon}})\geq 1$ for every $\varepsilon\in(0,1]$ and every $c^{(\pm)}\neq 0$.
  5. 5.
    Physical connection ($\varepsilon=1$): The bound holds for the physical connection $\nabla^{C}=\nabla^{C_{1}}$ as a direct consequence of the previous step.
This strategy rigorously applies the PT cohomological framework to the specific Kaluza–Klein geometry of the SOC BEC model by establishing that the essential topological constraint—encoded in $r^{\sharp}_{\varepsilon}=1$ for $\varepsilon>0$—translates into a geometric lower bound on off-diagonal holonomy through direct computation of the mixed-input curvature in the model’s parameter space.
5 Topological Lower Bound and Its Physical Consequences
With the geometric construction of Section 2 and the cohomological data of Section 3 at hand, we now synthesize these results to obtain the main physical consequence: a topological lower bound on the off-diagonal holonomy of the synthetic gauge connection in the SOC BEC model.
5.1 Statement of the bound
The central result is expressed by the following inequality, which follows from the uniform maximal rank of the obstruction map proved in Section 4:
$\dim\mathfrak{hol}^{\mathrm{off}}(\nabla^{C})\;\geq\;r^{\sharp}\;=\;\min(r-\dim\mathcal{K},1),$ (5.1)
where $r=\operatorname{rank}_{\mathbb{R}}([\omega]_{\mathrm{mixed}})$ is the mixed tensor rank of the torsion class, and $\mathcal{K}$ is the obstruction kernel. For the physical Kaluza–Klein metric $g_{M}$, we have established that $r^{\sharp}=1$.
5.2 Main theorem: topological lower bound for SOC BECs
Theorem 5.1 (Topological lower bound for SOC BECs).
Let $M=T_{\mathrm{BZ}}^{2}\times S_{\phi_{+}}^{1}\times S_{\phi_{-}}^{1}$ be the parameter space equipped with the Kaluza–Klein metric $g_{M}$ (Definition 2.5) and the metric connection $\nabla^{C}$ with torsion $T=F^{(+)}\wedge\Theta^{(+)}+F^{(-)}\wedge\Theta^{(-)}$. Under Assumption 2.3 (constant Berry curvatures), the off-diagonal holonomy algebra of $\nabla^{C}$ satisfies
$\dim\mathfrak{hol}^{\mathrm{off}}(\nabla^{C})\;\geq\;1\qquad\text{for every point }p\in M.$ (5.2)
Proof.
The proof follows from the uniform deformation analysis of Section 4:
  1. 1.
    Topological and geometric invariants: The mixed tensor rank is $r=1$ (Theorem 3.6). For $\varepsilon>0$ the obstruction kernel vanishes, $\mathcal{K}_{\varepsilon}=0$ (Lemma 2.16), giving reduced rank $r^{\sharp}_{\varepsilon}=1$.
  2. 2.
    Uniform bound for all $\varepsilon>0$: Theorem 4.5 proves that the obstruction map $\overline{\Psi}_{p}^{(\varepsilon)}$ has maximal rank $r^{\sharp}=1$ for every $\varepsilon\in(0,1]$ and every point $p$. By Ambrose–Singer, this yields $\dim\mathfrak{hol}^{\mathrm{off}}_{p}(\nabla^{C_{\varepsilon}})\geq 1$ uniformly (Corollary 4.6).
  3. 3.
    Physical endpoint: Since $\varepsilon=1$ belongs to the interval $(0,1]$, the bound holds directly for the physical connection $\nabla^{C}$.
Remark 5.2 (Methodological scope).
The adaptation strategy employed in this work does not rely on specific features of the SOC BEC model beyond the constant curvature assumption (Assumption 2.3). The key ingredients are:
  1. 1.
    a smooth deformation family $\{g_{\varepsilon}\}_{\varepsilon\in[0,1]}$ interpolating between a Riemannian product metric ($\varepsilon=0$) and the physical metric ($\varepsilon=1$);
  2. 2.
    explicit computation of the parallel-form strata $P^{k}(M_{i},g_{\varepsilon})$ for the deformed metrics;
  3. 3.
    verification that the obstruction kernel $K_{\varepsilon}$ has reduced dimension for $\varepsilon>0$ compared to the product limit.
This framework applies whenever the total space admits a fibered structure with computable cohomology and parallel forms. Natural candidates include Kaluza–Klein reductions in supergravity, principal bundles with connection in Yang–Mills theory, and synthetic gauge field systems in photonics and condensed matter physics.
5.3 Physical interpretation: non-removable off-diagonal curvature
Inequality (5.2) means that the synthetic gauge curvature cannot be made block-diagonal with respect to the splitting $T_{p}M=\mathcal{H}_{p}\oplus\mathcal{V}_{p}$. At least one independent off-diagonal curvature operator persists, mixing momentum with the phase directions $\phi_{\pm}$. This represents a topological obstruction to simultaneous flattening of the Berry curvature along both phase circles.
5.4 Persistence of the bound when the total Chern number vanishes
A crucial feature is that the bound is insensitive to the total first Chern number $c_{1}^{\mathrm{tot}}=c_{1}^{(+)}+c_{1}^{(-)}$.
Corollary 5.3 (Bound independent of total Chern number).
If $c_{1}^{(+)}\neq 0$ and $c_{1}^{(-)}=-c_{1}^{(+)}$ (so $c_{1}^{\mathrm{tot}}=0$), Theorem 5.1 still gives
$\dim\mathfrak{hol}^{\mathrm{off}}(\nabla^{C})\geq 1.$
Thus the topological obstruction persists even when the net Berry flux vanishes.
Proof.
As long as both $F^{(\pm)}\neq 0$, the mixed tensor rank is $r=1$ and $\mathcal{K}_{\varepsilon}=0$ (Lemma 2.16). The bound follows directly from Theorem 5.1. ∎
5.5 Robustness and experimental signatures
The bound depends only on topological data ($\omega]$ and parallel‑form strata), making it robust against perturbations. This suggests interferometric protocols to detect correlated Berry phases in the two U(1) sectors.
Proposition 5.4 (Experimental signature).
Under Assumption 2.3, no smooth gauge transformation can simultaneously eliminate the Berry phases $\Phi_{B}^{+}$ and $\Phi_{B}^{-}$ accumulated along any non‑contractible loop $\gamma\subset T^{2}_{\mathrm{BZ}}$. The persistent off-diagonal holonomy guarantees at least one non‑zero component of $(\Phi_{B}^{+}(\gamma),\Phi_{B}^{-}(\gamma))$ for every $\gamma$.
5.6 Summary
The inequality $\dim\mathfrak{hol}^{\mathrm{off}}(\nabla^{C})\geq 1$ provides a cohomologically defined, locally non‑trivial obstruction to complete Berry curvature removal in SOC BECs, valid even when global Chern invariants vanish. This extends the Chern‑number paradigm to a finer classification based on mixed cohomology classes.
6 Physical Consequences and Experimental Signatures
The topological lower bound established in Theorem 5.1 has direct physical consequences for the synthetic gauge structure of spin-orbit-coupled BECs. These consequences manifest as irreducible geometric couplings between momentum and phase degrees of freedom, with observable signatures in interferometric measurements.
6.1 Geometric obstruction to gauge flattening
The inequality $\dim\mathfrak{hol}^{\mathrm{off}}(\nabla^{C})\geq 1$ implies that the curvature tensor of the synthetic gauge connection cannot be made block-diagonal with respect to the splitting $T_{p}M=\mathcal{H}_{p}\oplus\mathcal{V}_{p}$. In more physical terms, no smooth gauge transformation or continuous deformation of the condensate order parameter can simultaneously eliminate the mixed components of the curvature that couple momentum to phase directions.
Definition 6.1 (Mixed curvature components).
Let $X\in\mathcal{H}_{p}$ be a horizontal vector (momentum direction) and $Z\in\mathcal{V}_{p}$ a vertical vector (phase direction). The mixed curvature components are defined as:
$\Omega(X,Z):=R^{C}(X,Z)\in\mathfrak{so}(T_{p}M),$
which map horizontal directions to vertical ones and vice versa.
Theorem 5.1 guarantees that at least one such mixed component is non-zero at every point $p\in M$. This represents a topological obstruction to flattening the synthetic gauge connection simultaneously along both phase circles $\phi_{+}$ and $\phi_{-}$.
6.2 Interpretation in terms of Berry curvature non-integrability
The mixed holonomy bound has a direct interpretation in terms of the non-integrability of the Berry connection. For a two-component SOC BEC, the synthetic gauge fields $A^{(\pm)}$ give rise to Berry curvatures $F^{(\pm)}=dA^{(\pm)}$. The topological lower bound implies that even if the net Chern number vanishes, the Berry connection cannot be simultaneously made flat in both $U(1)$ sectors.
Corollary 6.2 (Non-flattenability of Berry curvature).
Under Assumption 2.3, suppose $c^{(+)}\neq 0$ and $c^{(-)}\neq 0$. Then there does not exist a smooth gauge transformation
$A^{(\pm)}\mapsto A^{(\pm)}+d\lambda^{(\pm)}$
that makes both curvatures vanish identically, i.e., such that $F^{(+)}=F^{(-)}=0$ everywhere on $T^{2}_{\mathrm{BZ}}$.
This result is stronger than the usual Chern number obstruction: the Chern number only prevents making the curvature exact, but here we show that even when the Chern numbers cancel ($c_{1}^{(+)}=-c_{1}^{(-)}$), the curvature cannot be made to vanish pointwise in both sectors simultaneously.
6.3 Experimental signatures via interferometry
The geometric obstruction predicted by Theorem 5.1 suggests concrete interferometric tests. Consider measuring the Berry phases accumulated along closed loops $\gamma\subset T^{2}_{\mathrm{BZ}}$ separately for the two phase sectors $\phi_{+}$ and $\phi_{-}$.
Definition 6.3 (Two-component Berry phase vector).
For a closed loop $\gamma$ in the Brillouin zone, define:
$\Phi_{B}(\gamma)=\begin{pmatrix}\Phi_{B}^{+}(\gamma)\\ \Phi_{B}^{-}(\gamma)\end{pmatrix}$,
where $\Phi_{B}^{\pm}(\gamma)=\oint_{\gamma}A^{(\pm)}$ are the Berry phases for the two $U(1)$ sectors.
If the synthetic gauge field were fully reducible (i.e., if the holonomy were block-diagonal), one could choose gauges where both $\Phi_{B}^{+}(\gamma)$ and $\Phi_{B}^{-}(\gamma)$ vanish for every loop $\gamma$. The bound $\dim\mathfrak{hol}^{\mathrm{off}}\geq 1$ forbids such simultaneous vanishing.
6.4 Comparison with Chern number physics
The mixed-rank invariant $r^{\sharp}$ provides a finer classification than the Chern number in several important respects:
Chern number Mixed rank invariant $r^{\sharp}$
Global invariant (integrated over BZ) Local structural invariant
Quantizes net vorticity Quantifies irreducible coupling
Vanishes when fluxes cancel Non-zero when individual fluxes are non-zero
Detects topological charge Detects topological entanglement
Table 1: Comparison between Chern number and mixed-rank invariant.
The key distinction is that the Chern number $c_{1}^{\mathrm{tot}}$ can vanish due to cancellation of local contributions, while $r^{\sharp}$ remains non-zero as long as both individual curvatures are non-vanishing. This makes $r^{\sharp}$ particularly valuable for detecting topological effects in compensated systems or synthetic vacuum configurations.
6.5 Robustness and experimental feasibility
The topological lower bound is robust against typical experimental perturbations for two reasons:
  1. 1.
    Topological invariance: The mixed tensor rank $r$ depends only on cohomology classes, which are stable under small deformations of the Hamiltonian.
  2. 2.
    Open condition: The non-parallelism condition $\nabla^{LC}T\neq 0$ is open in the space of metrics and connections.
Current experimental capabilities in cold-atom systems are sufficient to test these predictions:
  • Independent phase control: Optical ”painting” techniques [8] allow independent manipulation of $\phi_{+}$ and $\phi_{-}$.
  • Berry phase measurement: Interferometric methods used to map Berry curvature in Hofstadter bands [2] can be adapted to track two phase sectors.
  • Momentum-space tomography: Time-of-flight measurements provide access to the Brillouin zone geometry.
A proposed experiment would:
  1. 1.
    Prepare a two-component SOC BEC in a toroidal trap.
  2. 2.
    Adiabatically transport the condensate along independent cycles $\gamma_{1},\gamma_{2}$ in the Brillouin zone.
  3. 3.
    Measure the accumulated phases $\Phi_{B}^{+}$ and $\Phi_{B}^{-}$ separately using interferometry.
  4. 4.
    Verify that no gauge transformation can simultaneously eliminate both phases.
6.6 Implications for topological quantum matter
The results presented here extend the classification of synthetic gauge fields beyond conventional topological invariants. The mixed-rank framework:
  1. 1.
    Reveals hidden topology: Detects topological obstructions invisible to Chern numbers.
  2. 2.
    Quantifies entanglement: The rank $r$ measures the irreducible coupling between different degrees of freedom.
  3. 3.
    Generalizes to multi-component systems: The approach naturally extends to systems with more than two internal states.
This suggests that mixed cohomology classes provide a powerful tool for characterizing the topological structure of synthetic gauge fields in engineered quantum systems, with potential applications beyond cold atoms to photonic systems, superconducting circuits, and other synthetic quantum platforms.
7 Illustrative Examples: From Chern Numbers to the Mixed Rank
The general lower bound $\dim\mathfrak{hol}^{\mathrm{off}}(\nabla^{C})\geq 1$ derived in Theorem 5.1 establishes a universal constraint for generic two-component SOC BECs. In this section, we illustrate its operational significance in two paradigmatic scenarios: first, the standard Rashba–Dresselhaus texture, where our framework recovers the phenomenology of Chern-number-induced obstructions; second, a configuration with vanishing net Chern flux, where the mixed-rank invariant $r^{\sharp}$ reveals topological features inaccessible to traditional global invariants.
7.1 Example 1: Rashba–Dresselhaus texture and the maximal rank $r=1$
Consider a two-dimensional, two-level Hamiltonian of the Rashba–Dresselhaus (RD) type:
$\hat{H}_{\mathrm{RD}}(k)\;=\;\frac{1}{2m}\bigl(\mathbf{p}-\mathbf{A}_{\mathrm{syn}}(k)\bigr)^{2}\;+\;\mathbf{B}_{\mathrm{syn}}(k)\cdot\boldsymbol{\sigma},\qquad k=(k_{x},k_{y})\in T^{2}_{\mathrm{BZ}},$
where $\mathbf{A}_{\mathrm{syn}}(k)$ and $\mathbf{B}_{\mathrm{syn}}(k)$ represent the synthetic vector potential and Zeeman field, respectively. For generic coupling parameters $\kappa$ and $\Omega$, the lower dressed band induces a spin texture map $\mathbf{n}:T^{2}_{\mathrm{BZ}}\to S^{2}$.
The associated eigenline bundle $\mathbb{L}\to T^{2}_{\mathrm{BZ}}$ has a first Chern number $c_{1}(\mathbb{L})$ equal to the degree of $\mathbf{n}$. For a minimal RD texture with degree $1$, the Berry curvature $F$ represents a non-trivial cohomology class $F]\in H^{2}(T^{2}_{\mathrm{BZ}};\mathbb{Z})$.
In the extended parameter space $M=T^{2}_{\mathrm{BZ}}\times S^{1}_{\phi_{+}}\times S^{1}_{\phi_{-}}$, the synthetic torsion class is
$\omega]=[F]\otimes[d\phi_{+}]\;+\;[F]\otimes[d\phi_{-}]\in H^{3}(M;\mathbb{R}).$
Since $\dim H^{2}(T^{2}_{\mathrm{BZ}})=1$, the two summands are linearly dependent, and the class factors as a single simple tensor:
$\omega]=[F]\otimes\bigl([d\phi_{+}]+[d\phi_{-}]\bigr).$
Hence, the mixed tensor rank is $r=1$. For the product metric $g_{0}$, the obstruction kernel has $\dim\mathcal{K}_{0}=1$, giving $r^{\sharp}_{0}=0$. However, for the physical Kaluza–Klein metric $g_{M}$ (or any $g_{\varepsilon}$ with $\varepsilon>0$), Theorem 3.14 shows $\mathcal{K}=0$, so $r^{\sharp}=1$. By the direct curvature analysis (Section 4) we obtain
$\dim\mathfrak{hol}^{\mathrm{off}}(\nabla^{C})\;\geq\;1.$
Remark 7.1 (Local vs. global structure).
When $c_{1}(\mathbb{L})\neq 0$, the underlying principal bundle is topologically non‑trivial. Nevertheless, our geometric analysis is performed in local trivializations, where $M$ is diffeomorphic to $T^{4}$ and the metric takes the form (2.1). The deformation argument and the pointwise injectivity of the obstruction map are local in nature and can be applied in each trivialization. Thus, the bound $\dim\mathfrak{hol}^{\mathrm{off}}\geq 1$ holds locally, detecting the irreducible coupling between momentum and phase directions even when the global bundle is non‑trivial.
7.2 Example 2: Topological persistence under vanishing net Chern flux
The power of the mixed-rank invariant becomes most apparent when global topological charges cancel. Consider a configuration where the two $U(1)$ sectors carry Berry curvatures with equal and opposite integrated fluxes:
$\frac{1}{2\pi}\int_{T_{\mathrm{BZ}}^{2}}F^{(+)}=+1,\qquad\frac{1}{2\pi}\int_{T_{\mathrm{BZ}}^{2}}F^{(-)}=-1.$ (7.1)
Then the total Chern number $c_{1}^{\mathrm{tot}}=c_{1}^{(+)}+c_{1}^{(-)}=0$. Conventional Chern-number diagnostics would classify this regime as topologically trivial, yet our bound reveals a residual topological obstruction.
Under Assumption 2.3 we have $F^{(\pm)}=c^{(\pm)}\operatorname{vol}_{\mathrm{BZ}}$ with $c^{(+)}=-c^{(-)}$. The cohomology classes satisfy $F^{(-)}]=-[F^{(+)}]$. Hence, the torsion class becomes
$\omega]=[F^{(+)}]\otimes[d\phi_{+}]+[F^{(-)}]\otimes[d\phi_{-}]=[F^{(+)}]\otimes\bigl([d\phi_{+}]-[d\phi_{-}]\bigr),$
which again has mixed rank $r=1$. Since $F^{(\pm)}\neq 0$, the obstruction kernel vanishes for the physical metric, giving $r^{\sharp}=1$. Theorem 5.1 therefore guarantees
$\dim\mathfrak{hol}^{\mathrm{off}}(\nabla^{C})\geq 1,$
even though $c_{1}^{\mathrm{tot}}=0$.
To see the bound concretely, choose the explicit connections
$A^{(+)}=\frac{1}{4\pi}(k_{x}dk_{y}-k_{y}dk_{x}),\qquad A^{(-)}=-A^{(+)},$
which satisfy (7.1) on a torus of period $2\pi$ in each direction. Write $c:=c^{(+)}=1/(2\pi)$, so that $c^{(-)}=-c$. The torsion 3-form in adapted gauge is
$T=c\,dk_{x}\wedge dk_{y}\wedge(d\phi_{+}-d\phi_{-}).$
Following the method of Appendix B, we evaluate the curvature on mixed inputs (one horizontal, one vertical). The torsion endomorphisms give, at leading order, the quadratic torsion brackets:
$\pi_{\mathrm{off}}\!\bigl(\tfrac{1}{4}[T_{e_{1}},T_{e_{3}}]\bigr)$ $:\;e_{1}\mapsto\tfrac{c^{2}}{4}(e_{3}-e_{4}),\quad e_{3}\mapsto-\tfrac{c^{2}}{4}\,e_{1},\quad e_{4}\mapsto\tfrac{c^{2}}{4}\,e_{1},$
$\pi_{\mathrm{off}}\!\bigl(\tfrac{1}{4}[T_{e_{2}},T_{e_{3}}]\bigr)$ $:\;e_{2}\mapsto\tfrac{c^{2}}{4}(e_{3}-e_{4}),\quad e_{3}\mapsto-\tfrac{c^{2}}{4}\,e_{2},\quad e_{4}\mapsto\tfrac{c^{2}}{4}\,e_{2}.$
These two off-diagonal operators are linearly independent in $\mathrm{Hom}(\mathcal{H}_{p},\mathcal{V}_{p})\oplus\mathrm{Hom}(\mathcal{V}_{p},\mathcal{H}_{p})$: the first involves $e_{1}$ while the second involves $e_{2}$. With $c=1/(2\pi)$, the coefficient is $c^{2}/4=1/(16\pi^{2})\neq 0$, confirming that in this specific geometry
$\dim\mathfrak{hol}^{\mathrm{off}}(\nabla^{C})\;\geq\;2\;>\;1\;=\;r^{\sharp}.$
Remark 7.2 (Sharpness of the bound).
The explicit computation confirms that the bound $\dim\mathfrak{hol}^{\mathrm{off}}\geq 1$ is sharp in the sense that it cannot be improved without additional assumptions: there exist geometries (like the product metric $g_{0}$) where the off-diagonal holonomy is trivial, while for generic Kaluza–Klein metrics the bound is non-trivial and often exceeded. In the example above, two linearly independent mixed-input curvature operators generate at least a 2-dimensional off-diagonal holonomy algebra, illustrating that $r^{\sharp}=1$ is a lower bound; particular geometries may exhibit a richer off-diagonal holonomy.
7.3 Comparison: Global invariants vs. mixed rank
The distinction between $c_{1}^{\mathrm{tot}}$ and $r^{\sharp}$ reflects a fundamental dichotomy between integrated and structural topology:
Chern number $c_{1}^{\mathrm{tot}}$ Mixed rank $r^{\sharp}$
Global aggregate of curvature Local algebraic measure of coupling
Quantifies net vorticity Quantifies irreducible entanglement
Can vanish by cancellation Non‑zero if each factor curvature is non‑zero
Detects topological charge Detects topological locking
Table 2: Comparison of topological invariants.
The mixed rank $r$ counts the minimal number of simple tensors needed to represent the cohomology class $\omega]$ in the Künneth decomposition. In our geometry, $r=1$ whenever both individual Berry curvature classes $F^{(\pm)}]$ are non‑zero (and hence proportional, because $\dim H^{2}(T^{2})=1$). Thus $r$ measures the irreducible geometric coupling between the Brillouin‑zone torus and the phase fibres. The bound $\dim\mathfrak{hol}^{\mathrm{off}}\geq r^{\sharp}$ therefore constrains the algebraic complexity of the curvature, ensuring that the synthetic gauge field remains “locked” in a non‑trivial configuration even when the net topological charge vanishes.
7.4 Summary
The Rashba–Dresselhaus example shows that the mixed‑rank framework is consistent with established Chern‑number results, while the $c_{1}^{\mathrm{tot}}=0$ example demonstrates that it provides a finer classification, detecting locally non‑removable curvature that conventional global invariants miss. This makes $r^{\sharp}$ a superior tool for analyzing the stability of synthetic gauge fields in complex SOC BEC architectures, especially in compensated regimes where global topological charges cancel but local geometric obstructions persist.
8 Beyond the Chern Number: Local Geometric Constraints and Prospective Experimental Protocols
While the Chern number $c_{1}$ quantifies the net vorticity through a global integral over the Brillouin zone, the lower bound on the off-diagonal holonomy dimension, $r^{\sharp}$, acts as a local structural constraint. It obstructs the simultaneous annihilation of Berry curvature components along independent directions within the product geometry. This phenomenon can be visualized as the Berry curvature being constrained to move along topological “rail-tracks” in the configuration space: even when the total topological charge is tuned to zero, specific directional components remain geometrically locked and cannot be erased by any smooth gauge transformation.
In the two-component SOC BEC manifold $M=T^{2}_{\mathrm{BZ}}\times S^{1}_{\phi_{+}}\times S^{1}_{\phi_{-}}$, the bound $r^{\sharp}=1$ is non-degenerate. Under Assumption 2.3, the constant, non-zero Berry curvatures break the integrability of the horizontal distribution, ensuring $\mathcal{K}=0$. Consequently, the bound holds directly for the physical Kaluza–Klein metric $g_{M}$. The curvature of any metric connection calibrated by the mixed class $\omega]$ is thus irreducible to a block-diagonal form; at least one independent off-diagonal operator persists, fundamentally mixing momentum dynamics with the phase evolution of $\phi_{+}$ and $\phi_{-}$.
This geometric prediction suggests concrete interferometric tests. Consider the simultaneous measurement of the Berry phases $\Phi_{B}^{+}(\gamma)$ and $\Phi_{B}^{-}(\gamma)$ accumulated along a closed loop $\gamma\subset T^{2}_{\mathrm{BZ}}$ by separately addressing the global and relative phase sectors. If the gauge field were fully reducible (i.e., if the holonomy were block-diagonal with respect to the product splitting), one could choose a gauge where both phases vanish for every $\gamma$. The bound $r^{\sharp}=1$ forbids such a simultaneous vanishing across all loops. For any pair of independent cycles $\{\gamma_{1},\gamma_{2}\}$ in $T^{2}_{\mathrm{BZ}}$, the two-component Berry-phase vector:
$\boldsymbol{\Phi}_{B}=\bigl(\Phi_{B}^{+}(\gamma_{1}),\Phi_{B}^{-}(\gamma_{1}),\Phi_{B}^{+}(\gamma_{2}),\Phi_{B}^{-}(\gamma_{2})\bigr)$
cannot be continuously deformed to zero while preserving the mixed cohomology class $\omega]$.
Such measurements are within the reach of current cold-atom experimental capabilities. Interferometric techniques used to map Berry curvature in Hofstadter bands [2] or to probe spin-texture topology in SOC systems [9] can be adapted to track the two phase degrees of freedom $\phi_{\pm}$ independently [5]. The required control can be implemented via optical “painting” [8], allowing for precise spatiotemporal manipulation of the condensate’s phase profile.
Observing that these Berry-phase components cannot be simultaneously flattened would provide the first direct experimental signature of the topological obstruction captured by $r^{\sharp}$. This demonstrates that the mixed-cohomology framework not only yields a rigorous mathematical lower bound but also provides a feasible protocol for detecting locally irremovable Berry curvature, effectively moving the study of quantum gases beyond the global Chern-number paradigm.
9 Conclusions
We have demonstrated that a two-component spin–orbit-coupled (SOC) BEC possesses an intrinsic geometric obstruction that prevents the simultaneous flattening of Berry curvature along the independent phase directions $\phi_{\pm}$. This obstruction persists even when the net global Chern flux vanishes, identifying a topological regime that eludes traditional integrated invariants.
The extended parameter space $M\cong T^{2}_{\mathrm{BZ}}\times S^{1}_{\phi_{+}}\times S^{1}_{\phi_{-}}$, equipped with the physical Kaluza–Klein metric $g_{M}$ induced by synthetic gauge fields, carries a metric connection $\nabla^{C}$ with totally skew-symmetric torsion. Under Assumption 2.3 (constant Berry curvatures), we have established that the mixed tensor rank of the torsion class is $r=1$.
The core of our proof lies in the resolution of the Kaluza–Klein kernel problem. By introducing a smooth deformation family of metrics $g_{\varepsilon}$ (Definition 2.5), we have shown that while the original PT bound is trivial for the product metric ($\varepsilon=0$), the obstruction kernel $\mathcal{K}_{\varepsilon}$ vanishes for any $\varepsilon>0$ due to the non-parallelism of the vertical coframe (Theorem 2.13). By applying Theorem 4.7, we transferred the non-trivial bound from the deformed cases to the physical endpoint $\varepsilon=1$. This yields the sharp inequality:
$\dim\mathfrak{hol}^{\mathrm{off}}(\nabla^{C})\;\geq\;1.$ (9.1)
This algebraic constraint implies that momentum degrees of freedom are fundamentally “locked” to the phase sectors, an irreducible coupling that cannot be removed by smooth gauge transformations. The persistence of the bound in regimes with zero total Chern number—as illustrated in Section 7—highlights that the mixed-rank invariant $r^{\sharp}$ operates beyond the Chern-number paradigm, detecting structural features of the gauge field that integrated invariants fail to resolve.
Experimentally, this framework suggests interferometric protocols to resolve independent Berry phases (Section 8), providing a direct signature of locally irremovable curvature. Conceptually, it establishes a new mechanism for geometric robustness in synthetic gauge systems, controlled by the algebraic structure of mixed cohomology components.
In conclusion, beyond the specific application to spin–orbit-coupled BECs, the deformation strategy developed here—reducing a Kaluza–Klein geometry to its product limit while tracking the evolution of the obstruction kernel—provides a general template for extending the Pigazzini-Toda lower bound [12] to fibered geometries beyond the Riemannian product setting. The essential requirement is that the parallel-form strata be computable for the physical metric, enabling explicit evaluation of the reduced rank $r^{\sharp}$. We expect this approach to be applicable to other physical systems where synthetic gauge fields induce non-trivial geometric phases, including photonic topological insulators, superconducting qubit arrays, and higher-dimensional Kaluza–Klein compactifications in string theory.
Acknowledgements
The authors gratefully acknowledge stimulating discussions with colleagues in geometry and cold-atom physics that helped shape the formulation of these results.
Appendix A Levi-Civita Connection of the Deformed Metric
In this appendix, we derive the connection 1-forms and the covariant derivatives for the deformation family $g_{\varepsilon}$ defined in (2.1). We work in a local coordinate system $(x,y,\phi_{+},\phi_{-})$ on $M$ such that $g_{\mathrm{BZ}}=dx^{2}+dy^{2}$ and $F^{(\pm)}=c^{(\pm)}dx\wedge dy$.
A.1 Orthonormal Coframe
For a fixed $\varepsilon\in[0,1]$, an orthonormal coframe $\{e^{a}\}$ for $g_{\varepsilon}$ is:
$e^{1}=dx,\quad e^{2}=dy,\quad e^{3}=d\phi_{+}+\varepsilon A^{(+)},\quad e^{4}=d\phi_{-}+\varepsilon A^{(-)}.$
The structure equations $de^{a}+\omega^{a}_{\;b}\wedge e^{b}=0$ (with $\omega_{ab}=-\omega_{ba}$) determine the connection 1-forms.
Taking the exterior derivative of $e^{3}$ and $e^{4}$:
$de^{3}=\varepsilon dA^{(+)}=\varepsilon c^{(+)}e^{1}\wedge e^{2},\quad de^{4}=\varepsilon dA^{(-)}=\varepsilon c^{(-)}e^{1}\wedge e^{2}.$
The non-zero connection 1-forms are:
$\omega_{31}=-\omega_{13}$ $=-\frac{\varepsilon c^{(+)}}{2}e^{2},\quad\omega_{32}=-\omega_{23}=\frac{\varepsilon c^{(+)}}{2}e^{1},$
$\omega_{41}=-\omega_{14}$ $=-\frac{\varepsilon c^{(-)}}{2}e^{2},\quad\omega_{42}=-\omega_{24}=\frac{\varepsilon c^{(-)}}{2}e^{1}.$
A.2 Covariant Derivatives of the Coframe
The covariant derivative of the vertical 1-forms is $\nabla^{LC_{\varepsilon}}e^{a}=-\omega^{a}_{\;b}\otimes e^{b}$.
For $e^{3}$:
$\nabla^{LC_{\varepsilon}}e^{3}=\frac{\varepsilon c^{(+)}}{2}(e^{2}\otimes e^{1}-e^{1}\otimes e^{2}).$ (A.1)
Similarly for $e^{4}$:
$\nabla^{LC_{\varepsilon}}e^{4}=\frac{\varepsilon c^{(-)}}{2}(e^{2}\otimes e^{1}-e^{1}\otimes e^{2}).$ (A.2)
Note that for $\varepsilon=0$ (product metric), $\nabla^{LC_{0}}e^{3}=\nabla^{LC_{0}}e^{4}=0$, confirming that vertical forms are parallel in the product case. For $\varepsilon>0$, the vertical coframe is no longer parallel.
A.3 Christoffel Symbols
A.3.1 Computation of $\Gamma^{\phi_{+}}_{xy}$
We now derive the explicit formula for the mixed Christoffel symbol using the connection 1-forms computed above.
Direct computation from connection 1-forms.
At point $p$ in adapted gauge, the Christoffel symbols are related to the connection 1-forms by:
$\Gamma^{\phi_{+}}_{xy}\big|_{p}=\Gamma^{3}_{12}\big|_{p}=\omega_{32}(\partial_{x})\big|_{p},$ (A.3)
where we use the coordinate identification $x^{1}=x,x^{2}=y,x^{3}=\phi_{+}$.
From the connection 1-forms derived above:
$\omega_{32}=\frac{\varepsilon c^{(+)}}{2}e^{1}=\frac{\varepsilon c^{(+)}}{2}dx.$ (A.4)
Evaluating on $\partial_{x}$:
$\omega_{32}(\partial_{x})\big|_{p}=\frac{\varepsilon c^{(+)}}{2}\,dx(\partial_{x})=\frac{\varepsilon c^{(+)}}{2}.$ (A.5)
Therefore:
$\Gamma^{\phi_{+}}_{xy}\big|_{p}=\frac{\varepsilon c^{(+)}}{2}.$ (A.6)
Antisymmetric part and torsion verification.
For the other mixed component:
$\Gamma^{\phi_{+}}_{yx}\big|_{p}=\omega_{32}(\partial_{y})\big|_{p}=\frac{\varepsilon c^{(+)}}{2}\,dx(\partial_{y})=0.$ (A.7)
The antisymmetric part (related to torsion) is:
$\Gamma^{\phi_{+}}_{xy}-\Gamma^{\phi_{+}}_{yx}=\frac{\varepsilon c^{(+)}}{2}-0=\frac{\varepsilon c^{(+)}}{2}.$ (A.8)
Note that the full torsion 3-form $T_{\varepsilon}$ evaluated gives:
$T_{\varepsilon}(\partial_{x},\partial_{y},\partial_{\phi_{+}})\big|_{p}=c^{(+)},$ (A.9)
which differs from the Christoffel antisymmetric part by a factor. This is because the connection $\nabla^{C}$ with torsion (Proposition 2.11) has the relationship:
$\nabla^{C}_{X}Y=\nabla^{LC}_{X}Y+\frac{1}{2}T(X,Y,\cdot)^{\sharp},$ (A.10)
where the factor $1/2$ and the metric raising of the index account for the difference. The Christoffel symbols computed here are those of the Levi-Civita connection $\nabla^{LC}$, which enter the curvature formula used in Theorem 4.5.
Final result.
The mixed Christoffel symbols are exact (no higher-order corrections):
$\Gamma^{\phi_{+}}_{xy}\big|_{p}=\frac{\varepsilon c^{(+)}}{2},$ (A.11)
$\Gamma^{\phi_{-}}_{xy}\big|_{p}=\frac{\varepsilon c^{(-)}}{2},$ (A.12)
with all other mixed symbols vanishing. These are exact because the connection 1-forms $\omega_{3a}$, $\omega_{4a}$ are linear in $\varepsilon$ with constant coefficients (under Assumption 2.3), and no higher-order terms arise in the structure equations.
These formulas provide the explicit Christoffel symbols required for the curvature computation in Theorem 4.5 and Appendix B.
Appendix B Curvature Computation for the SOC BEC Model
We establish the non-vanishing of off-diagonal curvature for the connection $\nabla^{C_{\varepsilon}}$, thereby justifying Theorem 4.5. The computation uses the adapted gauge of Appendix A and exploits the homogeneity of the model under Assumption 2.3.
B.1 Notation and curvature decomposition
At a point $p\in M$ in adapted gauge, the orthonormal frame is
$e_{1}=\partial_{x},\quad e_{2}=\partial_{y}\qquad(\text{horizontal, }\mathcal{H}_{p}),\qquad e_{3}=\partial_{\phi_{+}},\quad e_{4}=\partial_{\phi_{-}}\qquad(\text{vertical, }\mathcal{V}_{p}).$
Definition B.1 (Off-diagonal endomorphism).
An endomorphism $A\in\mathfrak{so}(T_{p}M)$ is called off-diagonal (with respect to the splitting $\mathcal{H}_{p}\oplus\mathcal{V}_{p}$) if $A(\mathcal{H}_{p})\subset\mathcal{V}_{p}$ and $A(\mathcal{V}_{p})\subset\mathcal{H}_{p}$. We write $\pi_{\mathrm{off}}(A)$ for the off-diagonal projection of $A$.
The curvature of $\nabla^{C_{\varepsilon}}=\nabla^{\mathrm{LC}(\varepsilon)}+\tfrac{1}{2}T_{\varepsilon}$ decomposes as
$R^{C_{\varepsilon}}(X,Y)\;=\;R^{\mathrm{LC}(\varepsilon)}(X,Y)\;+\;\tfrac{1}{2}\bigl[(\nabla^{\mathrm{LC}(\varepsilon)}_{X}T_{\varepsilon})(Y,\cdot)-(\nabla^{\mathrm{LC}(\varepsilon)}_{Y}T_{\varepsilon})(X,\cdot)\bigr]^{\sharp}\;+\;\tfrac{1}{4}[T_{X},T_{Y}],$ (B.1)
where $T_{X}(Z):=T_{\varepsilon}(X,Z,\cdot)^{\sharp}$ and $T_{X},T_{Y}]:=T_{X}\circ T_{Y}-T_{Y}\circ T_{X}$.
To produce an off-diagonal holonomy generator we evaluate the curvature on mixed inputs: one horizontal, one vertical. Concretely, we compute
$\pi_{\mathrm{off}}\!\bigl(R^{C_{\varepsilon}}(e_{1},e_{3})\bigr)\;\in\;\mathrm{Hom}(\mathcal{H}_{p},\mathcal{V}_{p})\oplus\mathrm{Hom}(\mathcal{V}_{p},\mathcal{H}_{p}).$
B.2 Homogeneity of the model
Proposition B.2 (Homogeneity).
Under Assumption 2.3, the curvature tensor $R^{C_{\varepsilon}}$ is translation-invariant on $M=T^{2}_{\mathrm{BZ}}\times S^{1}_{\phi_{+}}\times S^{1}_{\phi_{-}}$. In particular, its components in the orthonormal frame $\{e_{1},e_{2},e_{3},e_{4}\}$ are independent of the point $p\in M$.
Proof.
The torsion 3-form at $p$ in adapted gauge reads $T_{\varepsilon}|_{p}=c^{(+)}e^{1}\!\wedge e^{2}\!\wedge e^{3}+c^{(-)}e^{1}\!\wedge e^{2}\!\wedge e^{4}$, with constant coefficients $c^{(\pm)}$. The connection 1-forms (Appendix A) are $\omega_{31}=-\tfrac{\varepsilon c^{(+)}}{2}e^{2}$, $\omega_{32}=\tfrac{\varepsilon c^{(+)}}{2}e^{1}$, $\omega_{41}=-\tfrac{\varepsilon c^{(-)}}{2}e^{2}$, $\omega_{42}=\tfrac{\varepsilon c^{(-)}}{2}e^{1}$, all with constant coefficients. Since the curvature 2-forms $\Omega^{a}{}_{b}=d\omega^{a}{}_{b}+\omega^{a}{}_{c}\wedge\omega^{c}{}_{b}$ and the torsion endomorphisms $T_{X}$ are computed from these constant data, every component of $R^{C_{\varepsilon}}$ is independent of $p$. ∎
Corollary B.3 (Pointwise implies global).
If $\pi_{\mathrm{off}}(R^{C_{\varepsilon}}(e_{i},e_{\alpha}))\neq 0$ at one point $p_{0}\in M$, then the same holds at every $p\in M$.
B.3 Torsion endomorphisms
The torsion at $p$ in adapted gauge is
$T_{\varepsilon}|_{p}\;=\;c^{(+)}\,e^{1}\!\wedge e^{2}\!\wedge e^{3}\;+\;c^{(-)}\,e^{1}\!\wedge e^{2}\!\wedge e^{4}.$
This expression is independent of $\varepsilon$ at $p$ (because $A^{(\pm)}|_{p}=0$ in adapted gauge).
Lemma B.4 (Torsion endomorphism table).
The nonzero values of the endomorphism $T_{e_{a}}(e_{b})=T(e_{a},e_{b},\cdot)^{\sharp}$ are:
$T_{e_{a}}(e_{b})$ $b=1$ $b=2$ $b=3$ $b=4$
$a=1$ 0 $c^{(+)}e_{3}+c^{(-)}e_{4}$ $c^{(+)}e_{2}$ $c^{(-)}e_{2}$
$a=2$ $c^{(+)}e_{3}-c^{(-)}e_{4}$ 0 $c^{(+)}e_{1}$ $c^{(-)}e_{1}$
$a=3$ $c^{(+)}e_{2}$ $c^{(+)}e_{1}$ 0 0
$a=4$ $c^{(-)}e_{2}$ $c^{(-)}e_{1}$ 0 0
Proof.
Direct computation from $T(e_{a},e_{b},e_{c})=c^{(+)}(e^{1}\!\wedge e^{2}\!\wedge e^{3})(e_{a},e_{b},e_{c})+c^{(-)}(e^{1}\!\wedge e^{2}\!\wedge e^{4})(e_{a},e_{b},e_{c})$ and the definition $T_{e_{a}}(e_{b})=\sum_{c}T(e_{a},e_{b},e_{c})\,e_{c}$. ∎
Remark B.5 (Type analysis).
Observe from the table:
  • $T_{e_{1}}(e_{2})=c^{(+)}e_{3}+c^{(-)}e_{4}\in\mathcal{V}$  ($\mathcal{H}\to\mathcal{V}$: off-diagonal).
  • $T_{e_{1}}(e_{3})=-c^{(+)}e_{2}\in\mathcal{H}$  ($\mathcal{V}\to\mathcal{H}$: off-diagonal).
  • $T_{e_{3}}(e_{1})=c^{(+)}e_{2}\in\mathcal{H}$  ($\mathcal{H}\to\mathcal{H}$: diagonal).
Thus the torsion itself already mixes horizontal and vertical directions. The quadratic torsion term $\tfrac{1}{4}[T_{X},T_{Y}]$ inherits this mixing.
B.4 Quadratic torsion for mixed inputs
We compute the commutator $T_{e_{1}},T_{e_{3}}]$ using the table of Lemma B.4.
Proposition B.6.
The commutator $T_{e_{1}},T_{e_{3}}]\in\mathfrak{gl}(T_{p}M)$ is given by:
$[T_{e_{1}},T_{e_{3}}](e_{1})$ $=(c^{(+)})^{2}\,e_{3}+c^{(+)}c^{(-)}\,e_{4},$ (B.2)
$[T_{e_{1}},T_{e_{3}}](e_{2})$ $=0,$ (B.3)
$[T_{e_{1}},T_{e_{3}}](e_{3})$ $=-(c^{(+)})^{2}\,e_{1},$ (B.4)
$[T_{e_{1}},T_{e_{3}}](e_{4})$ $=-c^{(+)}c^{(-)}\,e_{1}.$ (B.5)
Proof.
We compute each entry using $T_{e_{1}},T_{e_{3}}](e_{c})=T_{e_{1}}(T_{e_{3}}(e_{c}))-T_{e_{3}}(T_{e_{1}}(e_{c}))$:
Action on $e_{1}$: $T_{e_{3}}(e_{1})=c^{(+)}e_{2}$. Then $T_{e_{1}}(c^{(+)}e_{2})=c^{(+)}(c^{(+)}e_{3}+c^{(-)}e_{4})$. Also $T_{e_{1}}(e_{1})=0$, so $T_{e_{3}}(T_{e_{1}}(e_{1}))=0$. Result: $(c^{(+)})^{2}e_{3}+c^{(+)}c^{(-)}e_{4}$.
Action on $e_{2}$: $T_{e_{3}}(e_{2})=-c^{(+)}e_{1}$. Then $T_{e_{1}}(-c^{(+)}e_{1})=0$. Also $T_{e_{1}}(e_{2})=c^{(+)}e_{3}+c^{(-)}e_{4}$, $T_{e_{3}}(c^{(+)}e_{3}+c^{(-)}e_{4})=0$. Result: 0.
Action on $e_{3}$: $T_{e_{3}}(e_{3})=0$, so the first term vanishes. $T_{e_{1}}(e_{3})=-c^{(+)}e_{2}$, $T_{e_{3}}(-c^{(+)}e_{2})=-c^{(+)}(-c^{(+)}e_{1})=(c^{(+)})^{2}e_{1}$. Result: $0-(c^{(+)})^{2}e_{1}=-(c^{(+)})^{2}e_{1}$.
Action on $e_{4}$: $T_{e_{3}}(e_{4})=0$, so the first term vanishes. $T_{e_{1}}(e_{4})=-c^{(-)}e_{2}$, $T_{e_{3}}(-c^{(-)}e_{2})=-c^{(-)}(-c^{(+)}e_{1})=c^{(+)}c^{(-)}e_{1}$. Result: $0-c^{(+)}c^{(-)}e_{1}=-c^{(+)}c^{(-)}e_{1}$. ∎
Corollary B.7 (Off-diagonal projection).
The off-diagonal part of $T_{e_{1}},T_{e_{3}}]$ is:
$\pi_{\mathrm{off}}\bigl([T_{e_{1}},T_{e_{3}}]\bigr)\;:\quad\begin{cases}e_{1}\;\mapsto\;(c^{(+)})^{2}\,e_{3}+c^{(+)}c^{(-)}\,e_{4}&(\mathcal{H}\to\mathcal{V}),\\[3.0pt] e_{2}\;\mapsto\;0,\\[3.0pt] e_{3}\;\mapsto\;-(c^{(+)})^{2}\,e_{1}&(\mathcal{V}\to\mathcal{H}),\\[3.0pt] e_{4}\;\mapsto\;-c^{(+)}c^{(-)}\,e_{1}&(\mathcal{V}\to\mathcal{H}).\end{cases}$ (B.6)
Every entry in (B.6) is already off-diagonal (no projection is lost), and the map is nonzero whenever $c^{(+)}\neq 0$.
B.5 Contribution of the remaining curvature terms
The curvature $R^{C_{\varepsilon}}(e_{1},e_{3})$ also receives contributions from $R^{\mathrm{LC}(\varepsilon)}$ and from the torsion derivative $\tfrac{1}{2}[(\nabla_{e_{1}}T)(e_{3},\cdot)-(\nabla_{e_{3}}T)(e_{1},\cdot)]^{\sharp}$. We now verify that these are of lower order.
Lemma B.8 (Levi-Civita curvature for mixed inputs).
For mixed inputs $(e_{1},e_{3})$, the Levi-Civita curvature vanishes to leading order:
$R^{\mathrm{LC}(\varepsilon)}(e_{1},e_{3})=0.$ (B.7)
Proof.
The curvature 2-forms computed in Appendix A are all proportional to $e^{1}\!\wedge e^{2}$:
$\Omega^{1}{}_{2}=-\tfrac{\varepsilon^{2}}{4}\bigl[(c^{(+)})^{2}+(c^{(-)})^{2}\bigr]\,e^{1}\!\wedge e^{2},$
and $\Omega^{3}{}_{1}=\Omega^{3}{}_{2}=\Omega^{4}{}_{1}=\Omega^{4}{}_{2}=\Omega^{3}{}_{4}=0$. Evaluating any $\Omega^{a}{}_{b}$ on $(e_{1},e_{3})$ gives zero, since $e^{1}\!\wedge e^{2}(e_{1},e_{3})=0$. ∎
Lemma B.9 (Torsion derivative for mixed inputs).
The torsion-derivative contribution to $R^{C_{\varepsilon}}(e_{1},e_{3})$ is of order $\mathcal{O}(\varepsilon)$:
$\tfrac{1}{2}\bigl[(\nabla_{e_{1}}T_{\varepsilon})(e_{3},\cdot)-(\nabla_{e_{3}}T_{\varepsilon})(e_{1},\cdot)\bigr]^{\sharp}=\mathcal{O}(\varepsilon).$ (B.8)
Proof.
The torsion $T_{\varepsilon}|_{p}=c^{(+)}e^{1}\!\wedge e^{2}\!\wedge e^{3}+c^{(-)}e^{1}\!\wedge e^{2}\!\wedge e^{4}$ has constant coefficients in adapted gauge. Its covariant derivative $(\nabla_{X}T)(Y,Z,W)$ involves Christoffel corrections applied to each argument. Since the only nonzero connection coefficients are of order $\varepsilon$ (specifically, $\omega_{3a}\sim\varepsilon c^{(+)}$ and $\omega_{4a}\sim\varepsilon c^{(-)}$), the covariant derivative $\nabla T$ is $\mathcal{O}(\varepsilon)$. ∎
B.6 Main result: non-vanishing off-diagonal curvature
Theorem B.10 (Off-diagonal curvature for mixed inputs).
For all $c^{(\pm)}\neq 0$ and $\varepsilon>0$, the off-diagonal curvature at every point $p\in M$ is nonzero:
$\pi_{\mathrm{off}}\!\bigl(R^{C_{\varepsilon}}(e_{1},e_{3})\bigr)\;\neq\;0\qquad\text{for all }p\in M.$ (B.9)
Proof.
We evaluate the three terms in the curvature decomposition (B.1) for the mixed inputs $X=e_{1}\in\mathcal{H}_{p}$, $Y=e_{3}\in\mathcal{V}_{p}$.
Term 1 (Levi-Civita): By Lemma B.8, $R^{\mathrm{LC}(\varepsilon)}(e_{1},e_{3})=0$.
Term 2 (Torsion derivative): By Lemma B.9, this is $\mathcal{O}(\varepsilon)$.
Term 3 (Quadratic torsion): By Proposition B.6, the off-diagonal projection (B.6) gives at $\mathcal{O}(1)$:
$\pi_{\mathrm{off}}\!\bigl(\tfrac{1}{4}[T_{e_{1}},T_{e_{3}}]\bigr)\;:\quad e_{1}\mapsto\tfrac{(c^{(+)})^{2}}{4}\,e_{3}+\tfrac{c^{(+)}c^{(-)}}{4}\,e_{4},\qquad e_{3}\mapsto-\tfrac{(c^{(+)})^{2}}{4}\,e_{1}.$ (B.10)
For $c^{(+)}\neq 0$, the map $e_{3}\mapsto-\tfrac{(c^{(+)})^{2}}{4}e_{1}$ is nonzero.
Combining: The off-diagonal part of the curvature is
$\pi_{\mathrm{off}}\!\bigl(R^{C_{\varepsilon}}(e_{1},e_{3})\bigr)\;=\;\pi_{\mathrm{off}}\!\bigl(\tfrac{1}{4}[T_{e_{1}},T_{e_{3}}]\bigr)\;+\;\mathcal{O}(\varepsilon).$
At $\varepsilon=0$ the leading term is already nonzero (it equals (B.10)). Since the leading term is independent of $\varepsilon$ and the correction is continuous and $\mathcal{O}(\varepsilon)$, for all $\varepsilon\in[0,1]$ the total remains nonzero.
Precisely: the matrix element $\langle e_{1},\pi_{\mathrm{off}}(R^{C_{\varepsilon}}(e_{1},e_{3}))\,e_{3}\rangle=-\tfrac{(c^{(+)})^{2}}{4}+\mathcal{O}(\varepsilon)$, which is strictly negative for all $\varepsilon$ sufficiently small, and by continuity and the explicit polynomial dependence on $\varepsilon$, remains nonzero for all $\varepsilon\in[0,1]$ (one verifies that the $\mathcal{O}(\varepsilon)$ correction cannot cancel the leading term for finite $\varepsilon$, since the correction is at most linear in $\varepsilon$ while the leading term is a fixed nonzero constant).
By homogeneity (Proposition B.2), the result holds at every point $p\in M$. ∎
Remark B.11 (Role of each curvature input type).
The computation reveals a crucial structural point: the off-diagonal holonomy generator arises from the curvature with mixed inputs $R^{C}(e_{1},e_{3})$ (one horizontal, one vertical), not from the horizontal-input curvature $R^{C}(e_{1},e_{2})$.
The curvature $R^{C}(e_{1},e_{2})$ restricted to $\mathcal{V}_{p}$ measures the $\mathfrak{so}(\mathcal{V})$-component of the holonomy (the diagonal part, which rotates fiber directions among themselves). This component is also nonzero but does not contribute to $\mathfrak{hol}^{\mathrm{off}}$.
In contrast, $R^{C}(e_{1},e_{3})$ evaluated on the basis vectors produces endomorphisms that map $\mathcal{H}\to\mathcal{V}$ and $\mathcal{V}\to\mathcal{H}$, which is precisely the $\mathrm{Hom}(\mathcal{H},\mathcal{V})\oplus\mathrm{Hom}(\mathcal{V},\mathcal{H})$ component of $\mathfrak{so}(T_{p}M)$—the off-diagonal holonomy algebra $\mathfrak{hol}^{\mathrm{off}}$ in the sense of [12].
B.7 Global off-diagonal holonomy bound
Corollary B.12 (Global off-diagonal holonomy bound).
For all $c^{(\pm)}\neq 0$ and $\varepsilon\in(0,1]$:
$\dim\mathfrak{hol}^{\mathrm{off}}_{p}(\nabla^{C_{\varepsilon}})\;\geq\;1\qquad\text{for every }p\in M.$ (B.11)
Proof.
By Theorem B.10, the off-diagonal curvature operator $\pi_{\mathrm{off}}(R^{C_{\varepsilon}}(e_{1},e_{3}))\neq 0$ at every point $p\in M$. By the Ambrose–Singer theorem, this operator belongs to $\mathfrak{hol}_{p}(\nabla^{C_{\varepsilon}})$. Its off-diagonal projection is nonzero by construction, so $\mathfrak{hol}^{\mathrm{off}}_{p}(\nabla^{C_{\varepsilon}})\neq\{0\}$, yielding $\dim\mathfrak{hol}^{\mathrm{off}}_{p}\geq 1$. ∎
B.8 Explicit off-diagonal curvature for all mixed inputs
For completeness, we record the leading-order off-diagonal projection for all four mixed-input curvatures.
Proposition B.13 (All mixed-input off-diagonal curvatures).
At leading order $\mathcal{O}(1)$, the off-diagonal projections of the quadratic torsion term for all mixed input pairs are:
$\pi_{\mathrm{off}}\!\bigl(\tfrac{1}{4}[T_{e_{1}},T_{e_{3}}]\bigr)$ $:\;e_{1}\mapsto\tfrac{(c^{(+)})^{2}}{4}e_{3}+\tfrac{c^{(+)}c^{(-)}}{4}e_{4},\quad e_{3}\mapsto-\tfrac{(c^{(+)})^{2}}{4}e_{1},\quad e_{4}\mapsto-\tfrac{c^{(+)}c^{(-)}}{4}e_{1},$ (B.12)
$\pi_{\mathrm{off}}\!\bigl(\tfrac{1}{4}[T_{e_{1}},T_{e_{4}}]\bigr)$ $:\;e_{1}\mapsto\tfrac{c^{(+)}c^{(-)}}{4}e_{3}+\tfrac{(c^{(-)})^{2}}{4}e_{4},\quad e_{3}\mapsto-\tfrac{c^{(+)}c^{(-)}}{4}e_{1},\quad e_{4}\mapsto-\tfrac{(c^{(-)})^{2}}{4}e_{1},$ (B.13)
$\pi_{\mathrm{off}}\!\bigl(\tfrac{1}{4}[T_{e_{2}},T_{e_{3}}]\bigr)$ $:\;e_{2}\mapsto\tfrac{(c^{(+)})^{2}}{4}e_{3}+\tfrac{c^{(+)}c^{(-)}}{4}e_{4},\quad e_{3}\mapsto-\tfrac{(c^{(+)})^{2}}{4}e_{2},\quad e_{4}\mapsto-\tfrac{c^{(+)}c^{(-)}}{4}e_{2},$ (B.14)
$\pi_{\mathrm{off}}\!\bigl(\tfrac{1}{4}[T_{e_{2}},T_{e_{4}}]\bigr)$ $:\;e_{2}\mapsto\tfrac{c^{(+)}c^{(-)}}{4}e_{3}+\tfrac{(c^{(-)})^{2}}{4}e_{4},\quad e_{3}\mapsto-\tfrac{c^{(+)}c^{(-)}}{4}e_{2},\quad e_{4}\mapsto-\tfrac{(c^{(-)})^{2}}{4}e_{2}.$ (B.15)
In particular:
  • If $c^{(+)}\neq 0$: the pairs $(e_{1},e_{3})$ and $(e_{2},e_{3})$ yield nonzero off-diagonal operators.
  • If $c^{(-)}\neq 0$: the pairs $(e_{1},e_{4})$ and $(e_{2},e_{4})$ yield nonzero off-diagonal operators.
  • For $c^{(+)}\neq 0$ and $c^{(-)}\neq 0$: all four pairs produce nonzero off-diagonal curvature.
Proof.
Analogous to the proof of Proposition B.6, using the torsion table of Lemma B.4. ∎
B.9 Physical interpretation
Remark B.14 (Experimental significance).
The homogeneity of the SOC BEC model has direct experimental consequences:
  1. 1.
    Uniform constraint: The off-diagonal curvature is nonzero at every point in the extended parameter space, not just on a dense open set. Any measurement of Berry-phase correlations will detect the irreducible mixing of momentum and phase degrees of freedom, regardless of the momentum-space location.
  2. 2.
    Robustness: The leading off-diagonal contribution (Theorem B.10) is of order $\mathcal{O}(1)$ in the deformation parameter $\varepsilon$, arising from the quadratic torsion term. This is structurally stable: it depends only on the constants $c^{(\pm)}\neq 0$ and cannot be removed by small perturbations.
  3. 3.
    Mixed-input origin: The off-diagonal holonomy arises from the curvature with mixed (horizontal–vertical) inputs, reflecting the physical coupling between momentum transport ($\mathcal{H}$) and phase evolution ($\mathcal{V}$).